### v4

# A Persistence-Driven Edit Distance for Trees with Abstract Weights

Abstract
In this work we define a novel edit distance for trees considered with some abstract weights on the edges. The metric is driven by the idea of considering trees as topological summaries in the context of persistence and topological data analysis. Several examples related to persistent sets are presented. The metric can be computed with a dynamical binary linear programming approach. This framework is applied and further studied in other works focused on merge trees, where the problems of stability and merge trees estimation are also assessed.
Keywords: Topological Data Analysis, Merge Trees, Reeb Graphs, Graph Edit Distance , Tree Edit Distance , Binary Optimization
1 Introduction
Graphs are widely used across scientific disciplines. Whenever a system is discretized, or when relations among points, statistical units, or abstract objects need to be formalized, graphs naturally emerge as a fundamental modeling tool.
This ubiquity has led to the development of a broad array of mathematical techniques for analyzing and comparing graphs, ranging from differential calculus on discrete structures and graph metrics, to specialized algebraic frameworks. Over time, this has given rise to a rich and diverse landscape of graph-theoretic methods.
Within this broader context, trees occupy a particularly important role. Their structural simplicity, combined with considerable expressive power, offers an appealing balance between interpretability and computational tractability. As a result, trees have become central in many areas of science. Two especially prominent examples are phylogenetic trees Billera et al. (2001); Felsenstein and Felenstein (2004); Garba et al. (2021); Lueg et al. (2024) and clustering dendrograms Murtagh and Contreras (2017); Xu and Tian (2015), which are closely related structures. Both describe how a set of elements evolves or merges under some notion of similarity or agglomeration, and both illustrate the versatility of trees as a means of encoding hierarchical information.
In this paper, we approach trees from the perspective of topological data analysis (TDA) (Edelsbrunner and Harer, 2008), interpreting them as carriers of topological information. Our goal is to develop an edit distance tailored to analyzing populations of tree-shaped topological summaries.
TDA is a relatively recent framework for data analysis that leverages tools from algebraic topology (Hatcher, 2000) to extract shape-related features from data at the level of individual statistical units.
A typical TDA pipeline begins with a topological space and constructs a nested sequence of subspaces—known as a filtration—from which topological information is extracted using functors such as simplicial homology (Hatcher, 2000). This standard approach has been generalized in several directions, including filtrations indexed by more general partially ordered sets (e.g., $\mathbb{R}^{n}$ for $n\geq 2$) (Botnan and Lesnick, 2022), and the use of other mathematical objects such as Laplacians (Mémoli et al., 2022), path-connected components (Curry et al., 2022, 2021; Pegoraro, 2024a; Cavinato et al., 2022), zigzag homology (Carlsson and De Silva, 2010), cosheaves (De Silva et al., 2016) among others.
Previous Works on Edit Distances for Trees in TDA
The topological perspective we assume in this work stems from persistent sets, as defined in Carlsson and Mémoli (2013), which are functors of the form $F:(P,\leq)\rightarrow\operatorname*{Top}\xrightarrow{\pi_{0}}\operatorname*{Sets}$ (with $(P,\leq)$ being a poset, $\operatorname*{Top}$ being the category of topological spaces and $\pi_{0}$ being the functor of path-connected components). The most typical example being $(P,\leq)$ equal to $\{f^{-1}((-\infty,t])\}_{t\in\mathbb{R}}$ for some $f:X\rightarrow\mathbb{R}$, which amounts to using merge trees of real valued functions (Morozov et al., 2013; Curry et al., 2021), while, if $(P,\leq)$ is the poset of open intervals in $\mathbb{R}$, up to some gluing conditions, we recover Reeb cosheaves (De Silva et al., 2016).
Relations between posets and graphs have been extensively studied: for instance it is well-known that any poset can be represented with a directed acyclic graph (DAG) (Aho et al., 1972). Thus, it is no surprise that a common way to represent instances of $F:(P,\leq)\rightarrow\operatorname*{Sets}$ is via some kind of graphs, like merge trees or Reeb graphs. Accordingly, the use of graphs (possibly labeled Yan et al. (2019)) as summaries of persistent sets has produced a series of original definitions and results (Bauer et al., 2014; De Silva et al., 2016; Di Fabio and Landi, 2016; Bauer et al., 2020; Morozov et al., 2013; Gasparovic et al., 2019; Pont et al., 2022; Sridharamurthy et al., 2020; Wetzels et al., 2022; Sridharamurthy and Natarajan, 2021; Cardona et al., 2021; Touli, 2020; Curry et al., 2022) which are all driven by these novel scenarios.
A significant number of these works focus on defining suitable metric structures for merge trees and Reeb graphs through variants of edit distances. In particular, (Di Fabio and Landi, 2016; Bauer et al., 2020) are focused on sup-norm kinds of distances, where the difference between (Reeb) graphs is determined by the cost of the biggest modification needed to turn one graph into another. While Pont et al. (2022); Sridharamurthy et al. (2020); Wetzels et al. (2022); Sridharamurthy and Natarajan (2021) consider edit distances for merge trees which are more closely related to classical tree edit distances Tai (1979), where the objective is to minimize the total cost of a sequence of modifications converting one merge tree into another.
In this work, we propose an edit distance for weighted trees that aligns more closely with the second family of approaches mentioned above, but it differs from previous methods in two key ways:
  • Stability Properties: in Pegoraro (2024a) and Pegoraro and Secchi (2024) it is shown that, when adapted to merge trees, the metric we define here produces an edit distance $d_{E}$ satisfying the following stability relation :
    $d_{I}(T,T^{\prime})\leq d_{E}(T,T^{\prime})\leq 2(\text{size}(T)+\text{size}(T^{\prime}))d_{I}(T,T^{\prime})$
    where $d_{I}$ is interleaving distance between merge trees (Morozov et al., 2013) and $\text{size}(T)$ is the number of edges in $T$. Which is analogous to the relation between the bottleneck distance ($d_{B}$) and the $1$-Wasserstein metric ($W_{1}$) between persistence diagrams (PDs):
    $d_{B}(D,D^{\prime})\leq W_{1}(D,D^{\prime})\leq(\#D+\#D^{\prime})d_{B}(D,D^{\prime}).$
    See Edelsbrunner and Harer (2010). Among the edit distances presented in Pont et al. (2022); Sridharamurthy et al. (2020); Wetzels et al. (2022); Sridharamurthy and Natarajan (2021), only the one in Wetzels et al. (2022) was shown—in Pegoraro (2024a)—to posses some stability properties. This stability result also clarifies the theoretical distinction between our approach and those in Bauer et al. (2014); Di Fabio and Landi (2016); Bauer et al. (2020), all of which yield distances that are bounded from above by $d_{I}$. Since $d_{B}$ is the universal metric between PDs and $d_{I}$ is the universal one between merge trees, by analogy, one can further characterize the difference between $d_{E}$ and Di Fabio and Landi (2016); Bauer et al. (2020), via the easier and well understood difference between $d_{B}$ and $W_{1}$.
  • Abstract Weights: we investigate generalizations in which trees are enriched not only with edge lengths, but with weights taking values in more general metric spaces. This allows topological summaries to capture and represent a wider variety of information.
Main Contributions
As anticipated, the main contribution of this work is the introduction of a novel edit distance designed to compare trees interpreted as topological summaries. Increasingly, such trees are enriched with additional information, typically associated with the vertices or edges of the graph, as seen in various recent works (Ushizima et al., 2012; Curry et al., 2022; Pegoraro, 2024b; Curry et al., 2023a). To remain as general as possible, we allow abstract weights to be assigned to the edges of the trees.
Given the wide range of settings that emerge from this enriched perspective (as discussed in Section 5 and reflected in the diverse works built on the framework we establish) this paper focuses on the following core contributions:
  1. 1.
    Defining a general edit distance for trees with abstract weights;
  2. 2.
    Formalizing the abstract properties that edge weights must satisfy for the metric to be well-defined;
  3. 3.
    Developing a computational framework for efficiently computing the metric;
  4. 4.
    Beginning to explore the variety of edge weights that can be used, discussing their potential applications and how can they support diverse data analysis scenarios. We leave a deeper investigation into the interpretability of the metric in specific scenarios and the study of its stability properties to other works.
The edit distance we define stems from classical edit distances (Levenshtein et al., 1966; Tai, 1979; Bille, 2005; Gao et al., 2010; Lerouge et al., 2016), but add some deep modifications, which can be intuitively motivated by the following example, represented in Figure 1.
(a) A function (orange) with its associated merge tree.
(b) Another function (brown) with its associated merge tree in black.
(c) Comparison bewtween the merge tree in Figure 1(b), and the tree obtained from the one in Figure 1(a), removing the edge $(a,b)$.
Figure 1: Plots related to the example presented in the Introduction.
Consider the following function: $f(x)=x\cdot\sin(x)$ for $x\in[0,2\pi]$ Figure 1(a). Then take the sublevel set filtration $F(t)=f^{-1}((-\infty,t])$ and the persistent set $\pi_{0}\circ F$. The information of this persistent set can be represented with a tree with four vertices and three edges, as in Figure 1(a), along with a real-valued height function defined on its vertices - with the root being at height +$\infty$+ ∞. We can associate to any edge its length, obtained as the difference in height between its endpoints. Suppose that we were to remove from the (merge) tree the edge $(a,b)$, we would obtain the red (merge) tree in Figure 1(c). Vertex $a$ is now a degree $2$ vertex which represents a merging between path connected components (i.e. a local maximum in the function), which is no more happening. On a function level, this can be visualized as removing the local minimum in $2\pi$, for instance restricting the function on $[0,3/2\pi]$ Figure 1(b). Usual edit distances would regard the red and black weighted trees in the green box in Figure 1(c) as very different trees, having a different number of edges and very different weights associated to the edges. As a consequence also the black trees in the green boxes in Figure 1(a) and in Figure 1(b) would be very far apart. In our case however, we would like our metric to be much less sensitive and, for instance, not to distinguish between the black and red trees in trees Figure 1(c). For this reason, differently from other edit distances, we add the possibility to remove the vertex $a$ and join the two adjacent edges.
Interconnected Works
The framework introduced in this work lays the foundation for a series of developments explored in separate studies: the edit distance we define is applied across different contexts, each accompanied by dedicated theoretical results that both justify the specific application and reinforce the foundational framework developed in this manuscript.
We summarize such relations as follows:
  • All the works we mention in the following points rely heavily on Section 4.2 and 2 for developing theoretic results, and on Algorithm 1 for applications and case studies.
  • Merge trees: this general framework is used by Pegoraro (2024a) to define a metric for merge trees, which can be computed with Algorithm 1. With the help of some results contained in Pegoraro and Secchi (2024), which is more focused on the statistical aspects of such metric, Pegoraro (2024a) shows that the approach we pursue here induces a metric with stability properties which are analogous to the ones of the $1$-Wasserstein distance between persistence diagrams. Pegoraro and Secchi (2024) and Cavinato et al. (2022) present real data applications to show the practical effectiveness of the metric.
  • Functions defined on merge trees: Pegoraro (2024b) exploits the present work, exploring the case where weights are functions defined on the edges of the trees, obtaining another stability result.
Outline
Section 2 introduces the idea of editable spaces, which are the spaces we want the weights of our graphs to live in. Their properties are exploited in Section 3 to define dendrograms. Section 4 develops the novel edit distance and proves that it is well defined for trees considered up to degree $2$ vertices. Section 5 presents a series of general examples regarding persistent sets. Section 6 obtains an algorithm to compute the distance via linear binary optimization. Section 7 concludes the paper by highlighting the pros and cons of the approach along with possible further developments, with a brief comparison with graph edit distances.
2 Editable Spaces
As outlined in the introduction, the goal of this manuscript is to introduce a novel metric structure for working with weighted trees, where the term weighted is interpreted broadly: edge weights are defined via a map that assigns each edge a value in a metric space. To define this metric meaningfully, we require the metric spaces to satisfy certain properties. These conditions are essential to guarantee well-behaved transformations when modifying trees.
Definition 1
A set $X$ is called editable if the following conditions are satisfied:
  • (P1)
    $(X,d_{X})$ is a metric space
  • (P2)
    $(X,\odot,0_{X})$ is a commutative monoid (that is $X$ has an associative operation $\odot$ with zero element $0_{X}$)
  • (P3)
    the map $d_{X}(0_{X},\cdot):X\rightarrow\mathbb{R}$ is a map of monoids between $(X,\odot)$ and $(\mathbb{R},+)$: $d_{X}(0_{X},x\odot y)=d_{X}(0_{X},x)+d_{X}(0_{X},y)$.
  • (P4)
    $d_{X}$ is $\odot$ invariant, that is: $d_{X}(x,y)=d_{X}(z\odot x,z\odot y)=d_{X}(x\odot z,y\odot z)$
Note that in property (P3), $d_{X}(x\odot y,0_{X})=d_{X}(x,0_{X})+d_{X}(y,0_{X})$, implies that $x\odot y\neq 0_{X}$. Moreover (P3)-(P4) imply that the points $0_{X}$, $x$, $y$ and $x\odot y$ form a rectangle which can be isometrically embedded in an Euclidean plane with the Manhattan geometry (that is, with the norm $\mid\mid\cdot\mid\mid_{1}$): $d_{X}(x,x\odot y)=d_{X}(0_{X},y)$, $d_{X}(y,x\odot y)=d_{X}(0_{X},x)$ and $d_{X}(x\odot y,0_{X})=d_{X}(0_{X},x)+d(0_{X},y)$.
Exploiting the monoidal structure of an editable space, we can turn it into a partially ordered set.
Proposition 1
An editable space $X$ admits a poset structure given by:
$x\leq y\text{ if and only if }\exists r\in X\text{ such that }x\odot r=y.$
Proof
  • Reflexivity: $x\odot 0=x$ and thus $x\leq x$.
  • Antisymmetry: suppose $x\leq y$ and $x\geq y$. Then $x\odot r=y$ and $y\odot q=x$ for some $r,q\in X$. But then:
    $d_{X}(y,0)=$ $d_{X}(x\odot r,0)=d_{X}(x,0)+d_{X}(r,0)$
    $=$ $d_{X}(y\odot q,0)+d_{X}(r,0)=d_{X}(y,0)+d_{X}(q,0)+d_{X}(r,0),$
    which implies $r=q=0$ and so $x=y$.
  • Transitivity follows from associativity of $\odot$.
If the editable space $X$ is contained in a group $V$, the partial order can be further characterized.
Corollary 1
Consider $X$ editable space. If $X\subset V$ with $V$ being a group and the inclusion being a map of monoids, then for every $x\leq y$ in $X$, there exist one and only one $r\in X$ such that $x+r=y$.
2.1 Examples of Editable Spaces
We give some examples of editable spaces.
2.1.1 Positive Real Numbers
The set $(\mathbb{R}_{\geq 0},+,\mid\cdot\mid)$ is an editable space, as well as its subsets which are monoids, like $\mathbb{N}$. This editable space can be used to recover the classical notion of (positively) weighted graphs.
2.1.2 $1$-Wasserstein Space of Persistence Diagrams
Persistence diagrams (PDs) are among the most used topological summaries in TDA. They are used to represent sequences of homology groups, often called persistence modules, arising by applying homology functors to filtrations of topological spaces. PDs are frequently used due to their interpretable nature, being finite set of points in the plane whose $x$ and $y$ coordinates - $y\geq x$ represent the birth and death of homological features along the filtration.
There are many possible ways to introduce persistence diagrams and Wasserstein metrics, like Edelsbrunner and Harer (2010), Skraba and Turner (2020), or Divol and Lacombe (2021). In particular, Edelsbrunner and Harer (2010) is suitable for interested readers not familiar with the topic. For notational convenience here we follow the approach in Bubenik and Elchesen (2022), which we use in the proof of an upcoming proposition.
Definition 2 (adapted from Bubenik and Elchesen (2022))
Define $D(\mathbb{R}^{2}_{\Delta})$ as the set of finite formal sums of elements in:
$\mathbb{R}^{2}_{\Delta}=\{(b,d)\in R^{2}\mid 0\leq b\leq d\}.$
Similarly $D(\Delta)$ is the set of formal sums of elements in $\Delta=\{(b,b)\in\mathbb{R}^{2}\mid 0\leq b\}$. In both sets of formal sums we have an operation given by the sum of formal sums, which is commutative and has the empty formal sum as zero element. Given $D_{1},D_{2}\in D(\mathbb{R}^{2}_{\Delta})$ we say that $D_{1}\sim D_{2}$ if there exist $D_{3},D_{4}\in D(\Delta)$ such that $D_{1}+D_{3}=D_{2}+D_{4}$. We collect the equivalence classes of such relations in $D(\mathbb{R}^{2}_{\Delta},\Delta)$ and call them persistence diagrams, and indicate such an equivalence class with the notation $[D]$. The sum operation and the zero element are coherently defined on the quotient.
As a consequence of 2 we can write every diagram $[D]$ as a sum of points in $\mathbb{R}^{2}_{\Delta}$ each with coefficient $1$. This “decomposition” is clearly non-unique as it varies between elements in the same equivalence class. With an abuse of notation we may say that $x\in[D]$, meaning that $x=(b,d)$ is an off-diagonal point of $\mathbb{R}^{2}_{\Delta}$ (i.e. $d>b$) appearing in such decomposition of an element (in fact, in any element) in the equivalence class of $D$.
Definition 3 (adapted from Bubenik and Elchesen (2022))
The $1$-Wasserstein metric between persistence diagrams is:
$W_{1}([D],[D^{\prime}])=\inf_{\gamma:\widetilde{D}\rightarrow\widetilde{D}^{\prime}}\parallel x-\gamma(x)\parallel_{1},$
where $\gamma$ is any bijection between some $\widetilde{D}$ such that $\widetilde{D}\sim D$, and some $\widetilde{D}^{\prime}$ such that $\widetilde{D}^{\prime}\sim D^{\prime}$. The infimum in the definition of $W_{1}([D],[D^{\prime}])$ is always attained by some $\gamma$, and thus, it is always a minimum.
Proposition 2
The space $(D(\mathbb{R}^{2}_{\Delta},\Delta),W_{1},+,0)$ is editable.
Remark 1
We assumed 3 as it makes 2 easier to prove, compared to other formulations. However, for the sake of simplicity, we drop the notation $[D]$ in the remaining of the manuscript referring to a diagram simply as $D$, chosing the representative with no points of the form $(b,b)$.
2.1.3 Functions in Editable Spaces
Consider an editable space $(X,d_{X},\odot_{X},0_{X})$ and a measure space $(M,\mu)$. Then, the space of functions:
$L_{1}(M,\mu,X):=\{f:M\rightarrow X\mid\int_{M}d_{X}(0_{X},f(p))d\mu(p)<\infty\}/\sim,$
with $\sim$ being the equivalence relation of functions up to zero measure subsets and $\odot_{W}$ being defined pointwise - is editable. In fact, the function $p\mapsto d_{X}(0,f(p))$ is always non negative, so if properties (P3) and (P4) hold for a fixed $p\in M$, then they hold also for integrals. Call $W:=L_{1}(M,X)$, we verify (P3) as follows:
[…]
$=$ $\int_{M}d_{X}(f(p),0)+d_{X}(g(p),0)d\mu(p)=d_{W}(f,0)+d_{W}(g,0).$
And similarly for (P4).
2.1.4 Finite Products of Editable Spaces
Consider two editable spaces $X$ and $X^{\prime}$, that is $(X,\odot,0_{X})$ and $(X^{\prime},\diamond,0_{X^{\prime}})$ satisfying properties (P1)-(P4). Then $(X\times X^{\prime},\star,(0_{X},0_{X^{\prime}}))$ is an editable space, with $\star$ being the component-wise operations $\odot$ and $\diamond$, and the metric $d$ on $X\times X^{\prime}$ being the (possibly weighted) sum of the component-wise metrics of $X$ and $X^{\prime}$. For instance, $\mathbb{R}^{n}_{\geq 0}$ with the $1$-norm is an editable space.
3 Dendrograms
Now we introduce the objects we will work with: trees, with a function associating to each edge a weight in a fixed editable space.
Definition 4
A tree structure $T$ is given by a connected rooted acyclic graph $(V_{T},E_{T})$. We indicate the root of the tree with $r_{T}$. The degree of a vertex $v\in V_{T}$ is the number of edges which have that vertex as one of the endpoints. Any vertex with an edge connecting it to the root is its child and the root is its parent: this is the first step of a recursion which defines the parent and children relation for all vertices in $V_{T}.$ The vertices with no children are called leaves and are collected in the set $L_{T}$. The relation $child<parent$ generates a partial order on $V_{T}$ which induces an orientation on the edges. The edges in $E_{T}$ are identified in the form of ordered pairs $(a,b)$ meaning $a<b$ and $a\rightarrow b$. A subtree of a vertex $v$, called $sub_{T}(v)$, is the tree structure whose set of vertices is $\{x\in V_{T}\mid x\leq v\}$.
A key fact is that given a tree structure $T$, identifying an edge $(v,v^{\prime})$ with its lower vertex $v$, gives a bijection between $V_{T}-\{r_{T}\}$ and $E_{T}$, that is $E_{T}\cong V_{T}-\{r_{T}\}$ as sets. Given this bijection, we often use $E_{T}$ to indicate the vertices $v\in V_{T}-\{r_{T}\}$, to simplify the notation.
To avoid overlapping some terminology with already existing notions of trees and for ease of notation, we refer to a weighted tree, with the weight function taking values in an editable space, by calling it dendrogram, as formalized by the following definition.
Definition 5
A tree structure $T$ with a weight function $\varphi_{T}:E_{T}\rightarrow X-\{0_{X}\}$, with $X$ editable space, is called dendrogram. The space of $X$-valued dendrograms is indicated as $(\mathcal{T},X)$.
3.1 Finite Posets, DAGs and Transitive Reductions
In this section, we justify our focus on tree-shaped objects by outlining, at a high level, the various contexts in which such graphs arise in TDA. This also yields a conceptual blueprint for potential extensions of the present work.
We start from the definition of a directed graph.
Definition 6
A directed graph is made by a pair $G=(V,E)$ with $V$ being the set of its vertices and $E\subseteq V\times V$ the set of its directed edges.
Consider a finite partially ordered set $(P,\leq)$. It is well known that one can associate to $P$ a unique (up to directed graph isomorphism) directed acyclic graph (DAG) $G_{P}$ so that $G_{P}$ is in bijection with the objects in $P$ and $(a,b)\in E_{G_{P}}$ if and only if $a\leq b$. Moreover, the graph $G_{P}$ can be simplified without losing ordering information. We know formalize these constructions.
Definition 7 (Aho et al. (1972))
A directed graph $G=(V,E)$ is called transitive if $(v,v^{\prime})\in E$, whenever there is a directed path from $v$ to $v^{\prime}$. For every directed graph $G=(V,E)$ we define its transitive closure as the graph $G^{T}:=(V,E_{G^{T}})$, with:
$E_{G^{T}}=\bigcap_{\{E^{\prime}\subset V\times V\mid E\subset E^{\prime},(V,E^{\prime})\text{ transitive graph}\}}E^{\prime}.$
Theorem 1 (adapted from Aho et al. (1972))
For every $G=(V,E)$ DAG there is a unique graph $G^{t}=(V,E_{G^{t}})$ such that:
  • $(G^{t})^{T}=G^{T}$;
  • for every $E^{\prime}\subsetneq E_{G^{t}}$, $(V,E^{\prime})^{T}\neq G^{T}$.
Such graph is called the transitive reduction of $G$ and it is given by:
$E_{G^{t}}:=\bigcap_{\{E^{\prime}\subset V\times V\mid(V,E^{\prime})^{T}=G^{T}\}}E^{\prime}.$
Given a finite poset $P$, we call $\mathcal{G}(P)$ the transitive reduction of the associated DAG.
3.2 From Persistent Sets to Dendrograms
We now establish a connection between functors, posets, and tree. To this end, we introduce a general procedure that applies beyond the specific case of functors $S:(\mathbb{R},\leq)\rightarrow\operatorname*{Sets}$, called persistent sets (Carlsson and Mémoli, 2013; Curry, 2018; Curry et al., 2022), which constitutes the primary focus of this work. Such objects are in fact of great interest in TDA and represent a (non-equivalent) alternative to persistence modules/diagrams when working with information collected from path-connected components of filtrations. A more thorough exploration of the broader scenarios is left for future research.
The objects we introduce in the following are represented in Figure 2.
Definition 8 (adapted from Curry et al. (2022))
Given a functor $S:P\rightarrow\operatorname*{Sets}$, with $P$ being a poset, the display poset of $S$ is defined as the set of pairs $D_{S}:=\{(p,s)\mid p\in P,s\in S(p)\}$. This is a poset with $(p,s)\leq(p^{\prime},s^{\prime})$ if and only if $p\leq p^{\prime}$ and $S(p\leq p^{\prime})(s)=s^{\prime}$. For every display poset $D_{S}$ we have a monotone increasing projection $h:D_{S}\rightarrow P$ defined by $(p,s)\mapsto p$.
The poset structure on $P$ and the one on $D_{S}$ are consistent with their respective graph representations, as made precise in the following proposition.
Proposition 3
The map $h:D_{S}\rightarrow P$ induces a graph homomorphism (a function between vertices that also sends edges to edges) between $\mathcal{G}(D_{S})$ and $\mathcal{G}(P)$.
Proof  By construction, any time we have $(p,s)\leq(p^{\prime},s^{\prime})\leq(p^{\prime\prime},s^{\prime\prime})$, this also implies $p\leq p^{\prime}\leq p^{\prime\prime}$. Viceversa, any time we have $(p,s)\leq(p^{\prime\prime},s^{\prime\prime})$ and $p\leq p^{\prime}\leq p^{\prime\prime}$, by functoriality there exists also $(p^{\prime},s^{\prime})$ such that $(p,s)\leq(p^{\prime},s^{\prime})\leq(p^{\prime\prime},s^{\prime\prime})$.
Now we restrict our attention to persistent sets.
Definition 9 ((Patel, 2018))
A constructible persistent set is a functor $S:(\mathbb{R},\leq)\rightarrow\operatorname*{Sets}$ such that there is a finite collection of real numbers $\{t_{1}<t_{2}<\ldots<t_{n}\}$ such that:
  • $S(t<t^{\prime})=\emptyset$ for all $t<t_{1}$;
  • for $t,t^{\prime}\in(t_{i},t_{i+1})$ or $t,t^{\prime}>t_{n}$, with $t<t^{\prime}$, then $S(t<t^{\prime})$ is bijective.
The set $\{t_{1}<t_{2}<\ldots<t_{n}\}$ is called critical set and $t_{i}$ are called critical values. If $S(t)$ is always a finite set, then $S$ is a finite persistent set.
Note that we can always consider a minimal set of critical values (Pegoraro, 2024a), which constitute the critical poset of the persistent set. We also add a last requirement for the persistent sets we want to work with.
Definition 10 (Pegoraro (2024a))
A regular persistent set $S$ is a finite constructible persistent set such that, for every critical value $t_{i}$ and for every $\varepsilon>0$ small enough, $S(t_{i}\leq t_{i}+\varepsilon)$ is a bijective map.
At this point we have the following immediate result, which we state without proof.
Proposition 4
Given two regular persistent sets $S$ and $S^{\prime}$, then $S\simeq S^{\prime}$ if and only if:
  • they have the same minimal set of critical values $C=\{t_{1},\ldots,t_{n}\}$
  • $S_{\mid C}\simeq S^{\prime}_{\mid C}$, with $S_{\mid C}:C\hookrightarrow\mathbb{R}\rightarrow\operatorname*{Sets}$.
Without going into the details of the construction of a merge tree, we just show that tree-structures are the natural discrete tool to work with persistent sets. See also 2.
Proposition 5
Consider a finite regular persistence set $S$ and a set of critical values $C$. If $S(t)$ has cardinality $1$ for $t$ big enough, $\mathcal{G}(D_{S_{\mid C}})$ is a tree.
Remark 2
In the proof of 5, we actually show that the graph $\mathcal{G}(D_{S_{\mid C}})$ is equivalent up to degree $2$ vertices (as a directed graph) to the merge tree of $S$, upon removing from the merge tree the edge that goes to infinity. In fact, proving the actual statement of 5 reduces to observing that, within the setting considered, $\mathcal{G}(D_{S_{\mid C}})$ is a rooted directed acyclic graph, with the $parent>child$ relationship being induced by the poset relationship on $D_{S_{\mid C}}$. Note that this proposition further justifies our idea of working up to degree $2$ vertices.
[…]
(b) The persistent set $\pi_{0}(X_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})$.
(c) The display poset $D_{\pi_{0}(X_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})}$.
[…]
Figure 2: An example of a filtration along with the persistent set of its path-connected components, the associated display poset and its graph representation. The colors are used throughout the plots to highlight the relations between the different objects.
4 Edits of Dendrograms
Now we present the main contributions of the paper.
The approach we follow is to define a distance inspired by Edit Distances (Tai, 1979), but with substantial differences in the edit operations. The philosophy of these distances is to allow certain modifications of the base object, called edits, each being associated to a cost, and to define the distance between two objects as the minimal cost that is needed to transform the first object into the second with a finite sequence of edits. In this way, up to properly setting up a set of edits, one can formalize the deformation of a dendrogram modifying the local information induced by the weight function defined on the edges and the tree structure itself. On top of that, edit distances frequently enjoy some decomposition properties which simplify the calculations (Hong et al., 2017), which are notoriously very heavy (Hein et al., 1995). We also point out that, in literature, tree are often considered with weights being real numbers and avoid modeling the weight space as we do with the definition of editable spaces.
Given a space of dendrograms $(\mathcal{T},X)$, with $(X,\odot,0_{X})$ editable space, and given a dendrogram $(T,\varphi)$, we define our edits as operators $\{(T,\varphi)\}\rightarrow(\mathcal{T},X)$. So, an edit operation is not defined on the whole $(\mathcal{T},X)$, but only on $(T,\varphi)$, and its image can be defined up to isomorphism. Moreover, given two edits $e_{i}:\{(T,\varphi)\}\rightarrow(\mathcal{T},X)$ and $e_{j}:e_{i}(\{(T,\varphi)\})\rightarrow(\mathcal{T},X)$, we can consider their composition $e_{i}\circ e_{j}$. Any finite composition of edits is referred to as an edit path.
The distance $d_{E}$ we introduce differs substantially from previously defined edit distances, as it is tailored specifically for comparing topological summaries—roughly meaning that topologically irrelevant points can be removed from a tree at no cost.
Recall that, via the identification $E_{T}\cong V_{T}-\{r_{T}\}$, the upcoming transformations can be described equivalently in terms of either edges or vertices.
  • Shrinking an edge means changing the weight value of the edge with a non zero element in $X$. The inverse of this transformation is the shrinking which restores the original edge weight.
  • Deleting an edge $(v_{1},v_{2})$ results into a new tree, with the same vertices apart from $v_{1}$ (the lower one), and with the parent of the deleted vertex which gains all of its children.
    The inverse of deletion is the insertion of an edge along with its child vertex. We can insert an edge at a vertex $v$ specifying the child of $v$ and its children (that can be either none or any portion of the children of $v$) and the weight of the edge.
  • Lastly, we can eliminate a degree two vertex $v$, that is, a parent with an only child, connecting the two adjacent edges which arrive and depart from $v.$ The weight of the resulting edge is the sum of the weights of the joined edges. This transformation is the ghosting of the vertex $v$. Its inverse transformation is called the splitting of an edge.
(a) Starting and target weighted trees, with highlighted the vertices involved in the edit path.
(b) Deletions of $v_{1},v_{2},v_{3}$, in red.
(c) Ghostings of $v_{4},v_{5}$, in yellow.
(d) Shrinkings of $v_{6}$ and $v_{7}$ to match the weight, respectively, of $w_{6}$ and $w_{7}$, in green.
(e) Splitting edges via $w_{4}$ and $w_{5}$, in yellow.
(f) Insertions of $w_{1},w_{2},w_{3}$, in red.
Figure 3: (b)$\rightarrow$(e) form an edit path made from the left weighted tree in Figure 3(a) to the right one. The edit path can be represented with a mapping - Section 4.2 - consisting of the pairs $(v_{i},\mathfrak{D})$ for all the red vertices in Figure 3(b), $(v_{i},\mathfrak{G})$ for all the yellow vertices in Figure 3(c), $(v_{i},w_{i})$ for all the vertices associated via the green color in Figure 3(d), $(\mathfrak{G},w_{i})$ for all the yellow vertices in Figure 3(c) and $(\mathfrak{D},w_{i})$ for all the red vertices in Figure 3(f).
See Figure 3 for an example of an edit path between two graphs whose weights are edges lenghts - Section 2.1.1.
Remark 3
Shrinking and deletions are classical edits for trees, with shrinking being usually referred to as relabeling. The ghosting edit (and its inverse) is completely unusual and we are not aware of any previous work employing it. Its definition is largely driven by the TDA perspective presented in the Introduction and it is fundamental for the stability results contained in Pegoraro and Secchi (2024) and Pegoraro (2024a). Even in the context of graph edit distances (GEDs) (Ambauen et al., 2003; Gao et al., 2010; Zeng et al., 2009; Lerouge et al., 2016; Serratosa, 2021), we are not aware of any existing work that adopts definitions similar to ours. While vertex splitting and merging in GEDs bear some resemblance to our notions of insertion and deletion, there no correspondence for splittings and ghostings as we define them.
We also give the following definition.
Definition 11
Dendrograms are equal up to degree $2$ vertices if they become isomorphic after applying a finite number of ghostings or splittings. We write $(T,\varphi_{T})\cong_{2}(T^{\prime},\varphi_{T^{\prime}})$.
11 induces an equivalence relation which identifies the set of dendrograms inside $(\mathcal{T},X)$ that we want to treat as equal. We call $(\mathcal{T}_{2},X)$ the space of equivalence classes of dendrograms in $(\mathcal{T},X)$, equal up to degree $2$ vertices.
4.1 Costs of Edit Operations
Now we associate to every edit a cost, that is, a length measure in the space $(\mathcal{T},X)$. In light of this interpretation, we will often use the words “length” and “cost” interchangeably when referring to an edit path.
The costs of the edit operations are defined as follows:
  • if, via shrinking, an edge goes from weight $x$ to weight $y$, then the cost of such operation is $d_{X}(x,y)$;
  • for any deletion/insertion of an edge with weight $x$, the cost is equal to $d_{X}(x,0_{X})$;
  • the cost of ghosting/splitting operations is always zero. This is coherent with the properties of editable spaces: consider a ghosting merging two adjacent edges separated by a degree two vertex, with $x$ being the weight of one edge, and $y$ the cost of the other. The total weight of the two edges is the same before and after the ghosting: $\mid d_{X}(x\odot y,0_{X})-d_{X}(x,0_{X})-d_{X}(y,0_{X})\mid=0$.
Definition 12
Given two dendrograms $T$ and $T^{\prime}$ in $(\mathcal{T},X)$, define:
  • $\Gamma(T,T^{\prime})$ as the set of all finite edit paths between $T$ and $T^{\prime}$;
  • $cost(\gamma)$ as the sum of the costs of the edits for any $\gamma\in\Gamma(T,T^{\prime})$;
  • the multigraph edit distance as:
    $d_{E}(T,T^{\prime})=\inf_{\gamma\in\Gamma(T,T^{\prime})}cost(\gamma)$
By definition the triangle inequality and symmetry must hold, but, up to now, this edit distance is intractable; one would have to search for all the possible finite edit paths which connect two dendrograms in order to find the minimal ones. On top of that, having an edit which is completely “for free”, it is not even obvious that $d_{E}(T,T^{\prime})>0$ for some dendrograms. However, it is clear that $d_{E}$ induces a pseudo-metric on classes of dendrograms up to degree two vertices.
4.2 Mappings
We introduce a combinatorial tool, called a mapping, which plays a central role in developing the theoretical properties of $d_{E}$ and in bringing it into the realm of computability. In particular, we prove that mappings can be used to parametrize a finite set of edit paths, with distinct and interpretable characteristics, which is guaranteed to contain at least an optimal edit path between two dendrograms.
Definition 13
A mapping between two dendrograms $T$ and $T^{\prime}$ is a set $M\subset(E_{T}\cup\{\mathfrak{D},\mathfrak{G}\})\times(E_{T^{\prime}}\cup\{\mathfrak{D},\mathfrak{G}\})$ satisfying:
  • (M1)
    consider the projection of the Cartesian product $(E_{T}\cup\{\mathfrak{D},\mathfrak{G}\})\times(E_{T^{\prime}}\cup\{\mathfrak{D},\mathfrak{G}\})\rightarrow(E_{T}\cup\{\mathfrak{D},\mathfrak{G}\})$; we can restrict this map to $M$ obtaining $\pi_{T}:M\rightarrow(E_{T}\cup\{\mathfrak{D},\mathfrak{G}\})$. The maps $\pi_{T}$ and $\pi_{T^{\prime}}$ are surjective on $E_{T}$ and $E_{T^{\prime}}$, i.e. $E_{T}\subset\operatorname*{Im}(\pi_{T})$ and $E_{T^{\prime}}\subset\operatorname*{Im}(\pi_{T^{\prime}})$;
  • (M2)
    $\pi_{T}$ and $\pi_{T^{\prime}}$ are injective on $M\cap(E_{T}\times E_{T^{\prime}})$;
  • (M3)
    given $(a,b)$ and $(c,d)\in M\cap(V_{T}\times V_{T^{\prime}})$, $a>c$, if and only if $b>d$;
  • (M4)
    if $(a,\mathfrak{G})\in M$ (or analogously $(\mathfrak{G},a)$), then after applying all deletions of the form $(v,\mathfrak{D})\in M$, the vertex $a$ becomes a degree $2$ vertex. In other words: let $child(a)=\{b_{1},..,b_{n}\}$. Then there is exactly one $i$ such that for all $j\neq i$, for all $v\in V_{sub(b_{j})}$, we have $(v,\mathfrak{D})\in M$; and there is one and only one $c$ such that $c=\max\{x<b_{i}\mid(x,y)\in M$ for any $y\in V_{T^{\prime}}\}$.
We call $\operatorname*{Mapp}(T,T^{\prime})$ the set of all mappings between $T$ and $T^{\prime}$.
We can always have a mapping made by:
$\{(e,\mathfrak{D})\mid e\in E_{T}\}\cup\{(\mathfrak{D},e)\mid e\in E_{T^{\prime}}\}.$
In the following, we will see that such a mapping corresponds to the edit path which deletes all the edges of one dendrogram and inserts all the edges of the second.
We may refer to edges which appear in the pairs in $M\cap(V_{T}\times V_{T^{\prime}})$ as the paired or matched edges/vertices.
Remark 4
In literature usually mappings are subsets of $E_{T}\times E_{T^{\prime}}$ and their properties are equivalent to (M2) and (M3) (see Hong et al. (2017); Sridharamurthy et al. (2020) and references therein). In this way, mappings coherently maps edges into edges, respecting the tree ordering, inducing in this way the needed relabeling operations - see also 3. The differences with our definition are caused by the introduction of the ghosting edits. Accordingly, the way in which mappings parametrize edit paths in the present work is completely novel and, for instance, it establishes some (partial) ordering between the edits. Note that the set of mappings we consider is bigger than in usual edit distances - thus computing the distance has a higher computational cost.
Every $M\in\operatorname*{Mapp}(T,T^{\prime})$ parametrizes a set of edit paths, with identical cost, as follows:
  • $\gamma_{d}^{T}$ is made by the deletions to be done on $T$, that is, the pairs $(v,\mathfrak{D})$, executed in any order. So we obtain $T^{M}_{d}=\gamma_{d}^{T}(T)$, which is well defined and does not depend on the order of the deletions. Similarly, we define $\gamma_{d}^{T^{\prime}}$ as a path made by the deletions to be done on $T^{\prime}$, that is, the pairs $(\mathfrak{D},w)$, executed in any order, and obtain $T^{\prime M}_{d}=\gamma_{d}^{T^{\prime}}(T^{\prime})$.
  • One then proceeds editing $T^{M}_{d}$ by ghosting all the vertices $(v,\mathfrak{G})$ in $M$, in any order, getting a path $\gamma^{T}_{g}$ and the dendrogram $T_{M}:=\gamma^{T}_{g}\circ\gamma^{T}_{d}(T)$. As before, we can do an analogous procedure on $T^{\prime M}_{d}$, ghosting all the vertices $(\mathfrak{G},w)$ in $M$, in any order, and building a path $\gamma^{T^{\prime}}_{g}$, along with the dendrogram $T^{\prime}_{M}:=\gamma^{T^{\prime}}_{g}\circ\gamma^{T^{\prime}}_{d}(T^{\prime})$.
  • Since all the remaining points in $M$ are paired, the two dendrograms $T^{\prime}_{M}$ and $T_{M}$ must be isomorphic as tree structures. This is guaranteed by the properties of $M$. So one can shrink $T_{M}$ onto $T^{\prime}_{M}$, and the composition of the shrinkings, executed in any order, gives an edit path $\gamma_{s}^{T}$.
By construction $\gamma_{s}^{T}\circ\gamma_{g}^{T}\circ\gamma_{d}^{T}(T)=T^{\prime}_{M}$, and $(\gamma_{d}^{T^{\prime}})^{-1}\circ(\gamma_{g}^{T^{\prime}})^{-1}\circ\gamma_{s}^{T}\circ\gamma_{g}^{T}\circ\gamma_{d}^{T}(T)=T^{\prime}$. Where the inverse of an edit path is thought as the composition of the inverses of the single edit operations, taken in the inverse order.
We call $\gamma_{M}$ the set of all possible edit paths of the form $(\gamma_{d}^{T^{\prime}})^{-1}\circ(\gamma_{g}^{T^{\prime}})^{-1}\circ\gamma_{s}^{T}\circ\gamma_{g}^{T}\circ\gamma_{d}^{T}$, obtained by changing the order in which the edit operations are executed inside $\gamma_{d}$, $\gamma_{g}$ and $\gamma_{s}$. Even if $\gamma_{M}$ is a set of paths, its cost is well defined:
$cost(M):=cost(\gamma_{M})=cost(\gamma_{d}^{T})+cost(\gamma_{s}^{T})+cost(\gamma_{d}^{T^{\prime}}).$
We prove that there exists always a mapping $M$ such that the paths in $\gamma_{M}$ are optimal.
Theorem 2 (Main Theorem)
Given two dendrograms $T$ and $T^{\prime}$, for every finite edit path $\gamma$, there is a mapping $M\in\operatorname*{Mapp}(T,T^{\prime})$ such that $cost(M)\leq cost(\gamma)$.
Two corollaries follow immediately from the facts that $\operatorname*{Mapp}(T,T\textquoteright)$ ’ ) is finite and that any mapping with cost zero must consist entirely of ghostings and splittings.
Corollary 2
For any two dendrograms $T$ and $T^{\prime}$, the set $\operatorname*{Mapp}(T,T^{\prime})$ is finite. Thus, we have the following well defined pseudo-metric between dendrograms:
$d_{E}(T,T^{\prime})=\inf\{cost(\gamma)\mid\gamma\in\Gamma(T,T^{\prime})\}=\min\{cost(M)\mid M\in\operatorname*{Mapp}(T,T^{\prime})\}$
which we will refer to as the edit distance between $T$ and $T^{\prime}$.
Corollary 3
Given $T$ and $T^{\prime}$ dendrograms, $d_{E}(T,T^{\prime})=0$ if and only if $T$ and $T^{\prime}$ are equal up to degree $2$ vertices. In other words $d_{E}$ is a metric for dendrograms considered up to degree $2$ vertices.
To conclude this section, we define a particular family of mappings which help us in further restricting the search space for optimal edit paths.
Definition 14
A mapping $M\in\operatorname*{Mapp}(T,T^{\prime})$ has maximal ghostings if the following hold: $(v,\mathfrak{G})\in M$ if, and only if, $v$ is of degree $2$ after the deletions in $T$ and, similarly, $(\mathfrak{G},w)\in M$ if, and only if, $w$ is of degree $2$ after the deletions in $T^{\prime}$.
A mapping $M\in\operatorname*{Mapp}(T,T^{\prime})$ has minimal deletions if the following hold: $(v,\mathfrak{D})\in M$ implies that neither $v$ nor $parent(v)$ are of degree $2$ after applying all the other deletions in $T$ and, similarly, $(\mathfrak{D},w)\in M$ implies that neither $w$ nor $parent(w)$ are of degree $2$ after applying all the other deletions in $T^{\prime}$.
We collect all mappings with maximal ghostings in the set $M_{\mathcal{G}}(T,T^{\prime})$ and paths with minimal deletions in the set $M_{\mathcal{D}}(T,T^{\prime})$. Lastly, we set: $M_{2}(T,T^{\prime}):=M_{\mathcal{G}}(T,T^{\prime})\bigcap M_{\mathcal{D}}(T,T^{\prime})$.
In other words with mappings in $M_{2}(T,T^{\prime})$ we are always eliminating all the degree $2$ vertices which arise from deletions and we are not deleting edges which we can shrink. The following lemma then applies.
Lemma 1
$\min\{cost(M)\mid M\in\operatorname*{Mapp}(T,T^{\prime})\}=\min\{cost(M)\mid M\in M_{2}(T,T^{\prime})\}.$
5 A Discussion on the Use of Different Editable Spaces
In this section we discuss some key examples which regard dendrograms with values in different editable spaces.
We stress again that this is to be intended as a roadmap, the outlining of a general idea and not as a formal and coherent description on how one can treat persistent sets via dendrograms. Coherently, we explicitly point out which questions and problems should be assessed by future works.
5.1 Function and Vector Valued Weights
[…]
  1. 1.
    diagram or Betti numbers enriched persistent sets, on the same line of thought of Curry et al. (2022, 2023b);
  2. 2.
    persistent sets enriched with the measure of sublevel sets.
[…]
We illustrate this with two examples: first we consider the datum of a point cloud a then a function defined on a open subset of $\mathbb{R}^{n}$.
Let $Y^{k}\subset\mathbb{R}^{n}$ be a finite set with $k$ points. A widely employed filtration is:
$Y_{t}^{k}:=\bigcup_{p\in Y^{k}}B_{t}(p)$
$B_{t}(p)$ being the metric ball centred in $p$, which is also called the offset filtration. Any point $D_{\pi_{0}(Y_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})}$ is a pair $(t,U)$, with $t\in\mathbb{R}$ and $U$ being a path connected component of $Y_{t}^{k}$. Then, one collect some topological information about $U$ and consider a vector-valued function $f$ with values in $X=\mathbb{R}^{N}_{\geq 0}$, for some $N<n$, defined by:
$f((t,U))=(\beta_{0}(U),\beta_{1}(U),\ldots,\beta_{N}(U)),$ (1)
with $\beta_{i}(U)=\dim\text{H}_{i}(U,\mathbb{R})$ being the $i$-th Betti number of $U$ clearly $\beta_{0}(U)=1$. For an example see Pegoraro (2024b). An alternative and richer way to capture shape-related information about $U$ is to consider the point cloud: $U_{k}:=Y^{k}\cap U$ and then build $\text{PD}_{i}(U_{k})$, the persistence diagram of the offset filtration of $U_{k}$ using $i$-dimensional homology. Consequently, one can define $f$ with values in $X=D(\mathbb{R}^{2}_{\Delta},\Delta)\times\ldots D(\mathbb{R}^{2}_{\Delta},\Delta)$, the space of persistence diagrams with the $1$-Wasserstein distance, defined by:
$f((t,U))=(\text{PD}_{0}(U_{k}),\ldots,\text{PD}_{N}(U_{k})).$
See Figure 4(a) and Figure 4(b) for a visual example. Observe how, at each value $t$, the merge tree gives a decomposition of the point cloud and the local shape of its clusters is captured by persistence diagrams.
Consider now a Lebesgue measurable function $g:W\rightarrow\mathbb{R}$, with $W$ open in $\mathbb{R}^{n}$. Build $Y_{t}=f^{-1}((-\infty,t])$, the sublevel set filtration of $f$. And consider $D_{\pi_{0}(Y_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})}$. This time another shape-related information that we can extract from a pair $(t,U)\in D_{\pi_{0}(Y_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})}$, is the Lebesgue measure of $U$. And thus define:
$f((t,U))=\mathcal{L}(U),$
with $\mathcal{L}(U)$ being the Lebesgue measure of $U$. For an example of situations that can be tackled with this approach see Figure 4(c) and Figure 4(d). Note that the representation of functional data obtained with the volume-enriched persistent sets, lies in between information captured by $L_{p}$ metrics and purely topological information: with persistent sets one would not be able to distinguish the two functions in Figure 4(c) and Figure 4(d) (they have isomorphic persistent sets), but with $L_{p}$ metrics one would not capture any of the shape similarities between the functions. We therefore believe that volume-enriched persistent sets can be a useful tool to employ when there are alignment issues with functional data i.e. all those situations where one may want to resort to some reparametrization of the domain - see, for instance, the Special Section on Time Warpings and Phase Variation on the Electronic Journal of Statistics, Vol 8 (2). One such example can be found in (Pegoraro, 2024b, Appendix,).
[…]
(a) A point cloud (left), with the thickening giving the off-set filtration at some particular value (dashed horizontal cyan line) and the display poset of (the persistent set of) the filtration (right). At the chosen value the filtration has three path connected components, shown in different colours, each with different shapes, characterized by the (qualitative) associated persistence diagram.
(b) Same situation as in Figure 4(a), but at a different filtration value. At the chosen value - bigger than in Figure 4(a), the filtration has two path connected components, shown in different colours, each with different shapes, characterized by the (qualitative) associated persistence diagram. At this resolution, the brown cluster shows a more persistent ringed shape.
(c) A real valued function defined on a square with four distinct local minima, associated to downward peaks with different shapes. The associated display poset (of the persistent set) accordingly shows four branches (black) and, in red, the (qualitative) volume of the associated downward peaks included in the sublevel sets, which increases with the value of the filtration.
(d) Same situation as in Figure 4(c), but for a different image. We clearly see that the four downward peaks have very different sizes, and that is reflected by the very different volume profiles attached to the branches of the merge tree. Note that the two images have the same merge trees.
Figure 4: Plots related to Section 5.1.
5.2 Problems to be Faced
In the previous section, we outlined ideas suggesting that different persistent sets can be represented by dendrograms equipped with weight functions taking values in various editable spaces. As noted at the beginning of Section 5, these ideas require formal validation before they can be reliably applied. In this section, we briefly discuss the main challenges associated with this proposed framework. Some of these challenges have been addressed in separate work, while others remain open and point to promising directions for future research.
There are three main issues that needs to be taken care of:
  1. 1.
    Discretization: Section 3.2 presents a procedure for turning a persistent set into a tree structure. This process depends crucially on the set of critical values $C$, and must be paired with the definition of an appropriate weight function, in order to obtain a dendrogram. Each of these steps—the discretization and the weight assignment—must be analyzed both individually and in combination. For example, one must verify whether the mapping from a persistent set to a dendrogram defines an injective transformation. This is not the case for the merge tree representation in Wetzels et al. (2022), while injectivity is ensured in Pegoraro (2024a, b).
  2. 2.
    Interpretability: the pipeline of representing a datum via an enriched persistent set and then with a dendrogram with values in an editable space, is in itself complex and needs to be carefully designed. Beyond the construction, it is also crucial to evaluate how well the edit distance $d_{E}$ captures variability between dendrograms. For instance, if the editable space is $\mathbb{R}_{\geq 0}$, it is clear that editing a positively weighted tree with our edits, always gives a positively weighted tree. However, when working with functions defined on display posets, Pegoraro (2024b) shows that only a subset of dendrograms corresponds to actual functions. Therefore, additional care is needed to ensure that $d_{E}$ meaningfully captures the variability between such functions.
  3. 3.
    Stability: closely related to interpretability is the issue of establishing stability guarantees specific to the chosen pipeline, akin to the results found in Pegoraro and Secchi (2024); Pegoraro (2024b). These results are essential to ensure that small perturbations in the data lead to correspondingly small changes in the associated dendrograms under $d_{E}$.
5.3 Stability Properties
In Section 5.2, we mention that establishing stability results is of pivotal importance to validate theoretically a pipeline involving dendrograms. Since some stability results have already been obtained in Pegoraro (2024a, b), we briefly describe them in order to give a qualitative idea of how the metrics $d_{E}$ behaves in different contexts.
In Pegoraro (2024a); Pegoraro and Secchi (2024) $d_{E}$ is extended (via some non-trivial steps) to merge trees. Building on such definition, as already mentioned in the introduction, it is therein shown that:
$d_{I}(T_{f},T_{g})\leq d_{E}(T_{f},T_{g})\leq 2\cdot\text{size}(T_{f})\cdot\text{size}(T_{g})\parallel f-g\parallel_{\infty}$ (2)
with:
  • $f,g$ being two tame (Chazal et al., 2016) real valued functions defined on the same path-connected topological space;
  • $T_{f}$ and $T_{g}$ are the merge trees associated to the sublevel set filtrations of $f$ and $g$;
  • $d_{I}$ is the interleaving distance between merge trees (Morozov et al., 2013);
  • $d_{E}$ is an adaptation of metric we define in this work, so that it induces a metric between merge trees;
  • $\text{size}(T_{f})$ being the cardinality of the edge set of $T_{f}$.
Equation 2 represents a stability property in line with our expectations: since our metric is a summation of all the costs of the local modifications needed to align two trees we have that the to control the distance between trees with the sup norm between functions, we need a constant that scales linearly with the size of the trees. The interleaving distance, instead, is focused on determining the maximal local modifications to turn one tree into the other and it is well known to be universally stable, in the sense that is the biggest distance which is always bounded by the sup norm between the functions. Moreover, we have already pointed out that the $1$-Wasserstein distance between persistence diagrams satisfies an analogous relation w.r.t. the bottleneck distances:
$d_{B}(D_{f},D_{g})\leq W_{1}(D_{f},D_{g})\leq(\#D_{f}+\#D_{g})\parallel f-g\parallel_{\infty}.$
As a consequence, our metric is better at picking up differences (analogously to Wasserstein distances w.r.t bottleneck distance for persistence diagrams). On the other hand, this also means that is it more sensitive to noise. Building on these considerations, in Pegoraro (2024a) it is argued that such stability properties may be preferable to universal stability properties, in which the operator from functions to topological representation is $1$-Lipschitz. Moreover, such comparison is framed in analogy to the bias-variance tradeoff in statistical modeling. We further observe that, since merge trees can be viewed as a special case of Reeb graphs, the same considerations we made for $d_{I}$ also apply to all the metrics discussed in Bollen (2022)—including those introduced in Bauer et al. (2014); Di Fabio and Landi (2016); Bauer et al. (2020)—due to their metric equivalence with the interleaving distance.
Lastly, Pegoraro (2024b) considers dendrograms enriched with Betti numbers of (path-connected components of) sublevel sets of functions, similarly to Equation 1. It is therein shown that:
$d_{E}(\varphi_{f}^{\Theta_{p}},\varphi_{g}^{\Theta_{p}})\leq(\#\text{Crit}(f)+\#\text{Crit}(g))2C\parallel f-g\parallel_{\infty},$
with:
  • $f,g$ being two tame (Chazal et al., 2016) real valued functions defined on the same path-connected topological space;
  • $\varphi_{f}^{\Theta_{p}}$ and $\varphi_{g}^{\Theta_{p}}$ are the dendrograms associated to (enriched) the persistent sets of the sublevel set filtrations of $f$ and $g$;
  • $d_{E}$ is an adaptation of metric we define in this work, so that it induces a metric between functions defined on display posets of persistent sets;
  • $C>0$ is an upper bound to the values of the Betti numbers;
  • $\text{Crit}(f)$ is the set of values where the homology of the sublevel sets of $f$ changes, in any dimension.
To conclude, we note that, if we consider the persistent homology of the sublevel set filtration of $f$ and $g$, in dimensions $i=0,\ldots,N$, and employ the persistence diagrams $D_{f}^{i}$ and $D_{g}^{i}$, we have:
$\sum_{i=0}^{N}W_{1}(D^{i}_{f},D^{i}_{g})\leq(\#\text{Crit}(f)+\#\text{Crit}(g))\parallel f-g\parallel_{\infty}.$
6 Computing the Edit Distance
In these last section we develop an algorithm to compute the edit distance for dendrograms.
6.1 Decomposition Properties
Following ideas found in Hong et al. (2017) and exploiting the properties of editable spaces, we now prove some theoretical results which allow us to recursively split up the problem of finding an optimal mapping between two dendrograms. Then, in Section 6.2, we formulate each of the split-up calculations as binary optimization problems, which are then aggregated recursively in a bottom-up fashion in Section 6.3·
Name $T_{2}$ the only representative without degree $2$ vertices inside the equivalence class of $T$. One can always suppose that a dendrogram is given without degree $2$ vertices. Thus, for notational convenience, from now on we suppose $T=T_{2}$ and $T^{\prime}=T^{\prime}_{2}$.
We consider some particular subsets of $E_{T}\times E_{T^{\prime}}$ which play a fundamental role in what follows. Recall that, using $E_{T}\cong V_{T}-\{r_{T}\}$, we can induce $\pi_{T}:E_{T}\times E_{T^{\prime}}\rightarrow V_{T}$.
A set $M^{*}\subset E_{T}\times E_{T^{\prime}}$ is in $\mathcal{C}^{*}(T,T^{\prime})$ if:
  • (A1)
    the points in $\pi_{T}(M^{*})$ form antichains in $V_{T}$ (and the same for $\pi_{T^{\prime}}(M^{*})$ in $V_{T^{\prime}}$), with respect to the partial order given by $parent>child$. This means that any two distinct vertices of $T$ (respectively of $T^{\prime}$) which appear in $M^{*}$ are incomparable with respect to “$>$”;
  • (A2)
    the projections $\pi_{T}:M^{*}\rightarrow V_{T}$ and $\pi_{T^{\prime}}:M^{*}\rightarrow V_{T^{\prime}}$ are injective.
Consider now $M^{*}\in\mathcal{C}^{*}(T,T^{\prime})$. Starting from such set of pairs we build a set of edits which form a “partial” mapping between $T$ and $T^{\prime}$: each pair $(x,y)\in M^{*}$ means that we do not care of what lies below $x\in V_{T}$ and $y\in V_{T^{\prime}}$ and we need to define edits only for the other vertices. The vertices below $x$ and $y$ will be taken care separately. See also Figure 5 for a visual example.
Loosely speaking, $M^{*}$ is used as a “dimensionality reduction tool”: instead of considering the problem of finding directly the optimal mapping between $T$ and $T^{\prime}$, we split up the problem in smaller subproblems, and put the pieces together using $M^{*}$. To formally do that, some other pieces of notation are needed.
Let $v\in E_{T}$. One can walk on the undirected graph of the tree-structure $T$ going towards any other vertex. For any $v\in E_{T}$, $\zeta_{v}$ is the shortest graph-path connecting $v$ to $r_{T}$. Note that this is the ordered set $\zeta_{v}=\{v^{\prime}\in V_{T}\mid v^{\prime}>v\}$. Similarly, denote with $\zeta_{x}^{x^{\prime}}$ the shortest path on the graph of $T$ connecting $x$ and $x^{\prime}$. Define the least common ancestor of a set of vertices $A$ as $LCA(A):=\min\{v\in V_{T}\mid v\geq A\}=\min\bigcap_{a\in A}\zeta_{a}$. In particular $\min\zeta_{x}\cap\zeta_{x^{\prime}}$ is the least common ancestor between $x$ and $x^{\prime}$: $LCA(x,x^{\prime})=\min\zeta_{x}\cap\zeta_{x^{\prime}}$.
By Property (A1), given $x\in V_{T}\cap\pi_{T}(M^{*})$, there exist a unique $\Omega_{M^{*}}(x)\notin\pi_{T}(M^{*})$ such that:
$\Omega_{M^{*}}(x)=\min\{LCA(x,x^{\prime})\mid x^{\prime}\in\pi_{T}(M^{*})\text{ and }x\neq x^{\prime}\}$
And the same holds for $y\in V_{T^{\prime}}\cap\pi_{T^{\prime}}(M^{*})$. For ease of notation we will often avoid explicit reference to $M^{*}$ and write directly $\Omega(x)$.
With these bits of notation, given $M^{*}\in\mathcal{C}^{*}(T,T^{\prime})$, we build the “partial” mapping $\alpha(M^{*})$: is a mapping that ignores all the vertices which lie below $x\in V_{T}$ and $y\in V_{T}$ if $(x,y)\in M^{*}$. Figure 5 may help in following the upcoming paragraph. Consider $v\in V_{T}$:
  1. 1.
    if $(v,w)\in M^{*}$, then $(v,w)\in\alpha(M^{*})$;
  2. 2.
    if there is not $x\in V_{T}$ such that $v<\Omega(x)$ or $v>\Omega(x)$, then $(v,\mathfrak{D})\in\alpha(M^{*})$;
  3. 3.
    if there is $x\in V_{T}$ such that $v>\Omega(x)$ then $(v,\mathfrak{D})\in\alpha(M^{*})$;
  4. 4.
    if there is $x\in V_{T}$ such that $v<\Omega(x)$:
    1. (a)
      if $v\in\zeta_{x}^{\Omega(x)}$ then $(v,\mathfrak{G})\in\alpha(M^{*})$
    2. (b)
      if $v<v_{i}$ for some $v_{i}\in\zeta_{x}^{\Omega(x)}=\{v_{0}<v_{1}<\ldots<v_{n}\}$ then $(v,\mathfrak{D})\in\alpha(M^{*})$;
    3. (c)
      if $v<x$ no edit is associated to $v$.
Remark 5
By Properties (A1) and (A2), the conditions used to build $\alpha(M^{*})$ are mutually exclusive. This means that each $v\in V_{T}$ satisfies one and only one of the above conditions and so $\alpha(M^{*})$ is well defined.
The idea behind $\alpha(M^{*})$ is that, for all pairs $(x,y)\in M^{*}$, we want to make the ghostings to turn the paths $\zeta_{x}^{\Omega(x)}$ and $\zeta_{y}^{\Omega(y)}$ respectively into the edges $(x,\Omega(x))$ and $(y,\Omega(y))$, and then shrink one in the other. As we already anticipated, $\alpha(M^{*})$ takes care of all the vertices in $V_{T}$ and $V_{T^{\prime}}$, a part from the sets $\cup_{(x,y)\in M^{*}}\{x^{\prime}\in E_{T}\mid x^{\prime}<x\}$ and $\cup_{(x,y)\in M^{*}}\{y^{\prime}\in E_{T^{\prime}}\mid y^{\prime}<y\}$. For this reason we say that $\alpha(M^{*})$ is a partial mapping.
We state this formally with the next proposition.
Proposition 6
Consider $T$ and $T^{\prime}$ and $M^{*}\in\mathcal{C}^{*}(T,T^{\prime})$. We obtain from such dendrograms, respectively, the dendrograms $\widetilde{T}$ and $\widetilde{T}^{\prime}$ by deleting all the vertices $\bigcup_{(x,y)\in M^{*}}\{x^{\prime}\in E_{T}\mid x^{\prime}<x\}$ and $\bigcup_{(x,y)\in M^{*}}\{y^{\prime}\in E_{T^{\prime}}\mid y^{\prime}<y\}$. The set $\alpha(M^{*})$ is a mapping in $M_{2}(\widetilde{T},\widetilde{T}^{\prime})$.
Now we have all the pieces we need to obtain the following key result.
Theorem 3 (Decomposition)
Given $T$, $T^{\prime}$ dendrograms:
$d_{E}(T,T^{\prime})=\min_{M^{*}\in\mathcal{C}^{*}(T,T^{\prime})}\sum_{(x,y)\in M^{*}}d_{E}(sub_{T}(x),sub_{T^{\prime}}(y))+cost(\alpha(M^{*})).$ (3)
Figure 5: Given two weighted trees $T$ (left) and $T^{\prime}$ (right) - which are the same of Figure 3(a) - we consider $T_{x}=sub_{T}(x)$ and $T^{\prime}_{y}=sub_{T^{\prime}}(y)$ and we use 3 to compute $d_{E}(T_{x},T^{\prime}_{y})$, as in the algorithm in Section 6.2. The set $M^{*}=\{(v,w),(v^{\prime},w^{\prime}),(v^{\prime\prime},w^{\prime\prime})\}$ satisfies (A1),(A2). The set $\alpha(M^{*})$ is made by the deletions/insertions indicated by the red edges, the ghostings/splittings indicated by the yellow and the shrinkings given by edges of the same color, different from red. To obtain $\widetilde{T}_{x}$ and $\widetilde{T}^{\prime}_{y}$ as in 6 all the black vertices covered by shaded regions must be deleted.
6.2 Dynamical Binary Linear Programming problems
We want to use 3 to write a dynamical, binary linear optimization algorithm to calculate $d_{E}$: by translating 3 into a Binary Linear Programming (BLP) problem, we obtain a single step in a bottom-up procedure.
6.2.1 Notation
We are given two dendrograms $T,T^{\prime}$ inside some dendrogram space $(\mathcal{T},X)$. Our objective is to write down Equation 3 as a function of some binary variables.
Consider $x\in V_{T}$ and $y\in V_{T^{\prime}}$. Along with keeping the notation defined in Section 6.1, define $T_{x}:=sub_{T}(x)$ and $T_{y}:=sub_{T^{\prime}}(y)$, $N_{x}:=\dim(T_{x})=\#E_{T}$ and $N_{y}:=\dim(T_{y})=\#E_{T^{\prime}}$. Moreover, given $v\in V_{T_{x}}$, the sequence $v_{0}=v<v_{1}<\ldots<r_{T}$ indicates the points in the path $\zeta_{v}$. Thus $v_{i}$ will be a vertex $v_{i}>v$. The same with $w\in V_{T_{y}}$
6.2.2 Relaxing the Optimization Problem
We would like to find $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$ minimizing Equation (3) for $T_{x}$ and $T_{y}$, but this is a difficult task. In fact, as evident in the construction of $\alpha(M^{*})$, a set $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$ has the role of pairing paths: if $(v,w)\in M^{*}$, then the paths $\zeta_{v}^{\Omega(v)}$ and $\zeta_{w}^{\Omega(w)}$ are shrunk one on the other by $\alpha(M^{*})$. However, the points $\Omega(v)$ and $\Omega(w)$ depend on the whole set $M^{*}$, and not simply on the pair $(v,w)$. Modeling such global dependence gives rise to non-linear relations between paired points, and so leading to a non linear cost function, in terms of points interactions, to be minimized. For this reason we “weaken” the last term in Equation (3), allowing also mappings different from $\alpha(M^{*})$ to be built from $M^{*}$. In other words we minimize over $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$ the following equation:
$\sum_{(v,w)\in M^{*}}d_{E}(sub_{T_{x}}(v),sub_{T_{y}}(w))+cost(\beta(M^{*}))$ (4)
where $\beta(M^{*})$ is such that:
  • $\beta(M^{*})\in M_{\mathcal{G}}(\widetilde{T}_{x},\widetilde{T}_{y})$ (using the notation established in 6, replacing $T$ and $T^{\prime}$ with $T_{x}$ and $T_{y}$ respectively);
  • the set of vertices paired by $\beta(M^{*})$ is exactly $M^{*}$: $M^{*}=\beta(M^{*})\bigcap V_{T_{x}}\times V_{T_{y}}$.
Since, by construction $M^{*}=\alpha(M^{*})\bigcap V_{T}\times V_{T^{\prime}}$ and by 6, $\alpha(M^{*})\in M_{2}(\widetilde{T}_{x},\widetilde{T}_{y})$, minimizing Equation (3) or Equation (4) gives the same results.
6.2.3 Setup and Variables
Suppose we already have $W_{xy}$ which is a $N_{x}\times N_{y}$ matrix such that $(W_{xy})_{v,w}=d_{E}(T_{v},T_{w})$ for all $v\in E_{T_{x}}$ and $w\in E_{T_{y}}$. Note that:
  • if $x$ and $y$ are leaves, $W_{xy}=0$.
  • if $v,w$ are vertices of $T_{x}$, $T_{y}$, then $W_{vw}$ is a submatrix of $W_{xy}$.
The function to be optimized is defined on the following set of binary variables: for every $v\in E_{T_{x}}$ and $w\in E_{T_{y}}$, for $v_{i}\in\zeta_{v}$, $v_{i}<r_{T_{x}}$, and $w_{j}\in\zeta_{w}$, $w_{j}<r_{T_{y}}$, take a binary variable $\delta^{v,w}_{i,j}$. We use $\delta$ to indicate the matrix of variables $(\delta^{v,w}_{i,j})_{v,w,i,j}$.
The mapping $\beta(M^{*})$ is built according to the variables $\delta^{v,w}_{i,j}$ with value $1$: we write a constrained optimization problem such that having $\delta^{v,w}_{i,j}=1$ means pairing the paths $\zeta_{v}^{v_{i+1}}$ (that is, the path which starts with $(v,v_{1})$ and ends with $(v_{i},v_{i+1})$) and $\zeta_{w}^{w_{j+1}}$, and shrinking one in the other in the induced mapping.
In order to pair and shrink the paths $\zeta_{v}^{v_{i+1}}=\{v=v_{0},v_{1},\ldots,v_{i+1}\}$ and $\zeta_{w}^{w_{j+1}}$ we need to collect some edits in the set $\beta(M^{*})$ adding the following edits:
  • all the points $v_{k}\in\zeta_{v}^{v_{i+1}}$ with $10<k<i+$ 1 are ghosted, that is $(v_{k},G)\in\beta(M^{*})$;
  • if $v^{\prime}<v_{k}$ for some $10<k<i+$ 1, then $(v^{\prime},D)\in\beta(M^{*})$;
  • if $v^{\prime}\geq v_{i+1}$ and $v^{\prime}\neq r_{T}$, then $(v^{\prime},D)\in\beta(M^{*})$
  • $(v,w)\in\beta(M^{*})$.
In the end, all edges which are not assigned an edit operation, are deleted. Note that analogous edits must be induced on vertices in $T_{y}$. Thus, the edit $(v,w)\in\beta(M^{*})$, along the edit paths induced by $\beta(M^{*})$, means: shrinking the edge $(v,v_{i+1})$ onto $(w,w_{j+1})$). Recall that, if $\delta^{v,w}_{i,j}=1$, we do not need to define edits for $sub_{T_{x}}(v)$ and $sub_{T_{y}}(w)$ since, by assumption, we already know $d_{E}(T_{v},T_{w})$.
6.2.4 Constraints
Clearly, the set of $\delta^{v,w}_{i,j}$ with value $1$ does not always produce $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$: for instance paths could be paired multiple times. To avoid such issues, we build a set of constraints for the variable $\delta$.
For each $v^{\prime}\in V_{T_{x}}$ we call $\Phi(v^{\prime}):=\{(v^{\prime\prime},i)\in V_{T}\times\mathbb{N}\mid v^{\prime}=v^{\prime\prime}_{i}\in\zeta_{v^{\prime\prime}}^{v^{\prime\prime}_{i+1}}\}$. In an analogous way we define $\Phi(w^{\prime})$ for $w^{\prime}\in V_{T_{y}}$. Call $\mathcal{K}$ the set of values of $\delta$ such that for each leaf $l$ in $V_{T_{x}}$:
$\sum_{v^{\prime}\in\zeta_{l}}\left(\sum_{(v^{\prime\prime},i)\in\Phi(v^{\prime})}\sum_{w,j}\delta^{v^{\prime\prime},w}_{i,j}\right)\leq 1$ (5)
and for each leaf $l^{\prime}$ in $V_{T_{y}}$:
$\sum_{w^{\prime}\in\zeta_{l^{\prime}}}\left(\sum_{(w^{\prime\prime},j)\in\Phi(w^{\prime})}\sum_{v,i}\delta^{v,w^{\prime\prime}}_{i,j}\right)\leq 1$ (6)
The following proposition clarifies the properties of any value of $\delta\in\mathcal{K}$.
Proposition 7
If $\delta\in\mathcal{K}$:
  • the pairs $(v,w)$ such that $\delta^{v,w}_{i,j}=1$ define a set $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$;
  • the edits induced by all $\delta^{v,w}_{i,j}=1$ give a mapping $\beta(M^{*})$ in $M_{``G^{\prime\prime}}(\widetilde{T}_{x},\widetilde{T}_{y})$. With $\widetilde{T}_{x},\widetilde{T}_{y}$ being obtained from $T_{x}$ and $T_{y}$ as in 6.
Remark 6
If for every $\delta^{v,w}_{i,j}=1$ we have $v_{i+1}=\Omega(v)$, then $\beta(M^{*})=\alpha(M^{*})$.
6.2.5 Objective Function
Having built a mapping $\beta(M^{*})$ using the binary variables, we want to define a cost functions which computes the cost of such mapping depending on $\delta$.
As before, consider $v\in E_{T_{x}}$ and interpret $\delta^{v,w}_{i,j}=1$ as pairing the paths $\zeta_{v}^{v_{i+1}}$ and $\zeta_{w}^{w_{j+1}}$; then $v$ is paired with some $w\in E_{T_{y}}$ if $C(v):=\sum_{i,w,j}\delta^{v,w}_{i,j}=1$ and is ghosted if $G(v):=\sum_{\{i,v^{\prime}\mid v\in\zeta_{v^{\prime}}^{v^{\prime}_{i}}\}}\sum_{w,j}\delta^{v^{\prime},w}_{i,j}=1$. The vertex $v$ is instead deleted if $D(v):=1-C(v)-G(v)=1$. We introduce also the following quantities, which correspond to the cost of shrinking $\zeta_{v}^{v_{i+1}}$ on $\zeta_{w}^{w_{j+1}}$:
$\Delta^{v,w}_{i,j}=d_{X}\left(\bigodot_{v^{\prime}\in\zeta_{v}^{v_{i}}}\varphi_{T_{x}}(v^{\prime}),\bigodot_{w^{\prime}\in\zeta_{w}^{w_{j}}}\varphi_{T_{y}}(w^{\prime})\right)$
Note that the above operations are taken inside the editable space $X$.
The function which computes the cost given by paired points is therefore:
$F^{C}(\delta):=\sum_{v,w,i,j}\Delta^{v,w}_{i,j}\cdot\delta^{v,w}_{i,j}$
The contribution of deleted points is: $F^{D}(\delta)-F^{-}(\delta)$, where
$F^{D}(\delta):=\sum_{v\in T_{x}}D(v)\cdot d(\varphi_{T_{x}}(v),0)+\sum_{w\in T_{y}}D(w)\cdot d(\varphi_{T_{y}}(w),0)$
and
$F^{-}(\delta):=\sum_{v\in T_{x}}C(v)\cdot\mid\mid sub_{T_{x}}(v)\mid\mid+\sum_{w\in T_{y}}C(w)\cdot\mid\mid sub_{T_{y}}(w)\mid\mid$
where the “norm” of a tree $T$ is $\mid\mid T\mid\mid=\sum_{e\in E_{T}}d_{X}(\varphi(e),0)$.
Finally, one must take into account the values of $d_{E}(T_{v},T_{w})$, whenever $v$ and $w$ are paired; this information is contained in $(W_{xy})_{v,w}$:
$F^{S}(\delta):=\sum_{v,w}(W_{xy})_{v,w}\cdot\left(\sum_{i,j}\delta^{v,w}_{i,j}\right)$
Proposition 8
With the notation previously introduced:
$d_{E}(T_{x},T_{y})=\min_{\delta\in\mathcal{K}}F^{C}(\delta)+F^{D}(\delta)-F^{-}(\delta)+F^{S}(\delta)$ (7)
Proof  The contribution of paired points is $F^{C}(\delta)$ and the contribution of deleted points is $F^{D}(\delta)-F^{-}(\delta)$.
The cost of $\beta(M^{*})$ is: $F^{\beta}(\delta):=F^{C}(\delta)+F^{D}(\delta)-F^{-}(\delta)$. Lastly, $F^{S}(\delta)$ takes into account the value of $d_{E}(T_{v},T_{w})$, if $v$ and $w$ are paired. By 3, combined with 7, the solution of the following optimization problem:
$\min_{\delta\in\mathcal{K}}F^{S}(\delta)+F^{\beta}(\delta)$ (8)
is equal to $d_{E}(T_{x},T_{y})$.
Remark 7
A solution to Problem Equation 7 exists because the minimization domain is finite and there are admissible values; it is not unique in general.
6.3 Bottom-Up Algorithm
In this section the results obtained in Section 6.1 and the formulation established in Section 6.2 are used to obtain the algorithm implemented to compute the metric $d_{E}$ between dendrograms. Some last pieces of notation are introduced in order to describe the “bottom-up” nature of the algorithm.
Given $x\in V_{T}$, define $\operatorname*{len}(x)$ to be the number of vertices in $\zeta_{x}$ and $\operatorname*{len}(T)=\max_{v\in V_{T}}\operatorname*{len}(v)$. Then, we set $\operatorname*{lvl}(x)=\operatorname*{len}(T)-\operatorname*{len}(x)$.
The key property is that: $\operatorname*{lvl}(x)>\operatorname*{lvl}(v)$ for any $v\in sub(x)$. Thus, if $W_{xy}$ is known for any $x\in\operatorname*{lvl}_{T}^{-1}(n)$ and $y\in\operatorname*{lvl}_{T^{\prime}}^{-1}(m)$, then for any $v$, $w$ in $V_{T}$, $V_{T^{\prime}}$ such that $\operatorname*{lvl}(v)<n$ and $\operatorname*{lvl}(w)<m$, $W_{vw}$ is known as well. With this notation we can write down Algorithm 1.
Result: $d_{E}(T,T^{\prime})$
initialization: $N=\operatorname*{len}(T)$, $M=\operatorname*{len}(T^{\prime})$, $n=m=0$;
while $n\leq N$ or $m\leq M$ do
       for $(x,y)\in V_{T}\times V_{T^{\prime}}$ such that $\operatorname*{lvl}(x)\leq n$ and $\operatorname*{lvl}(y)\leq m$ do
             Calculate $(W_{r_{T}r_{T^{\prime}}})_{x,y}$ solving Problem (7);
       end for
      $n=n+1$; $m=m+1$;
end while
return $(W_{r_{T}r_{T^{\prime}}})_{r_{T},r_{T^{\prime}}}$
Algorithm 1 Bottom-Up Algorithm.
An hands-on example of some iterations of Algorithm 1 can be found in Section A.
6.4 Computational Cost
We end up with a result to analyze the performances of Algorithm 1 in the case of dendrograms with binary tree structures.
Proposition 9 (Computational Cost)
Let $T$ and $T^{\prime}$ be two dendrograms with full binary tree structures with $\dim(T)=\#E_{T}=N$ and $\dim(T^{\prime})=M$.
Then $d_{E}(T,T^{\prime})$ can be computed by solving $O(N\cdot M)$ BLP problems with $O(N\cdot\log(N)\cdot M\cdot\log(M))$ variables and $O(\log_{2}(M)+\log_{2}(N))$ constraints.
Proof  In a full binary tree structure, at each level $l$ we have $2^{l}$ vertices. Let $L=\operatorname*{len}(T)$ and $L^{\prime}=\operatorname*{len}(T^{\prime})$. We have that, for any vertex $v\in\operatorname*{lvl}(l)^{-1}$, the cardinality of the path from $v$ to any of the leaves in $sub_{T}(v)$ is ${L-l}$ and the number of leaves in $sub_{T}(v)$ is $2^{L-l}$.
So, given $v\in V_{T}$ at level $l$ and $w\in V_{T^{\prime}}$ at level $l^{\prime}$, to calculate $d_{E}(sub_{T}(v),sub_{T^{\prime}}(w))$ (having already $W_{vw}$) we need to solve a binary linear problem with $2^{L-l}\cdot(L-l)\cdot 2^{L^{\prime}-l^{\prime}}\cdot(L^{\prime}-l^{\prime})$ variables and $2^{L-l}+2^{L^{\prime}-l^{\prime}}$ linear constraints.
Thus, to calculate $d_{E}(T,T^{\prime})$, we need to solve $(2^{L+1}-1)\cdot(2^{L^{\prime}+1}-1)$ linear binary optimization problems, each with equal or less than $2^{L}\cdot L\cdot 2^{L^{\prime}}\cdot L^{\prime}$ variables and equal or less that $2^{L}+2^{L^{\prime}}$ constraints. Substituting $L=\log_{2}(N)$ and $L^{\prime}=\log_{2}(M)$ in these equations gives the result.
Note that binary dendrograms are dense (with respect to $d_{E}$) in any dendrogram space as long as for any $\varepsilon>0$, there is $x\in(X,\odot,0_{X})$ such that $d_{X}(x,0)<\varepsilon$.
Remark 8
We end the manuscript with a remark about the computational cost of $d_{E}$ as described in 9. We find this result surprising to some extent. If we consider a recent implementation of the classical edit distance between unlabeled and unordered trees obtained with BLP (Hong et al., 2017), then the complexity of the two metrics is very similar: the classical edit distance can be computed by solving $O(N\cdot M)$ BLP problems with $O(N\cdot M)$ variables and $O(\log_{2}(M)+\log_{2}(N))$ constraints - $O(N+M)$ if we count also the constraints restricting the integer variables to $\{0,1\}$, as the authors of Hong et al. (2017) do. Thus, the introduction of the ghosting and splitting edits, with all the complexities they carry, just increases the cost by two $\log$ factors in the number of variables. Moreover, the condition used in Zhang (1996) to make the classical edit distance tractable (that is, item (2) in Section 3.1. of Zhang (1996)) is compatible with (M1)-(M4) and thus a constrained version of $d_{E}$ is well defined. We believe that it could be computed with a polynomial time algorithm. We leave this investigation to future works.
7 Conclusions
We develop a novel framework to work with tree that arise as topological summaries in the context of topological data analysis, possibly enriched with some abstract weights as a result of different pipelines. These kind of summaries are increasingly frequent and require some particular care to be properly handled by a metric. We define an edit distance employing some novel edit operations, proving that such definition induces a metric up to a certain equivalence relation between weighted trees.
By exploiting the ordering properties of trees, we produce a binary linear programming algorithm to compute such metric. The modifications that we make w.r.t. the classical edit distance between unlabeled trees make it only marginally more expensive, namely by some log factors. For this reason we argue that polynomial time approximations should be available also for $d_{E}$. Still, the computational procedure we describe here is used in separate applied works, as mentioned in the introduction. These additional modifications are essential for the stability properties assessed in Pegoraro and Secchi (2024); Pegoraro (2024b), justifying the higher computational cost w.r.t. other edit distances for merge trees, which, instead, are unstable.
The generality of the work opens up many possible research directions. Some of which are already studied in other works, while others still await for investigations:
  • we think that the properties of the editable spaces can be relaxed; however, the algorithm presented in this manuscript may need to be adapted to the properties of the chosen weight space;
  • we would like to extend this edit distance outside tree-shaped graphs, encompassing Reeb graphs (Biasotti et al., 2008; De Silva et al., 2016) and their combinatorial approximation, namely mapper graphs (Singh et al., 2007) are studied (Carriere et al., 2018; Carriere and Oudot, 2018), perhaps following the decomposition presented in Stefanou (2020). We also mention that works on GEDs like (Ambauen et al., 2003; Lerouge et al., 2016) could be of great help in developing such generalization;
  • as motivated by 8 we believe that polynomial time approximations can be obtained for computing $d_{E}$ between dendrograms;
  • further applications and case studies exploiting the generality of the weights we consider on the edges should be considered and explored.
Acknowledgments
This work was carried out as part of my PhD Thesis, under the supervision of Professor Piercesare Secchi. I also acknowledge the support of the Wallenberg AI, Autonomous Systems and Software Program (WASP), and of the SciLifeLab and Wallenberg National Program for Data-Driven Life Science (DDLS), which fund the project: Topological Data Analysis of Functional Genome to Find Covariation Signatures.
Outline of the Appendix
Section A contains an hands-on example of some iterations of the algorithm presented in Section 6.3. Lastly, Section B contains the proofs of the results in the paper.
A Algorithmic Example
Here we present in details the first steps of the Algorithm 1, used to calculate the distance between two merge trees.
We consider the following pair of weighted trees. Let $(T,w_{T})$ be the tree given by: $V_{T}=\{a,b,c,d,r_{T}\}$, $E_{T}=\{(a,d),(b,d),(d,r_{T}),(c,r_{T})\}$ and $w_{T}(a)=w_{T}(b)=w_{T}(d)=1$, $w_{T}(c)=5$; the tree $(T^{\prime},w_{T^{\prime}})$ instead, is defined by: $V_{T^{\prime}}=\{a^{\prime},b^{\prime},c^{\prime},d^{\prime},r_{T^{\prime}}\}$, $E_{T^{\prime}}=\{(a^{\prime},d^{\prime}),(b^{\prime},d^{\prime}),(d^{\prime},r_{T^{\prime}}),(c^{\prime},r_{T^{\prime}})\}$ and $w_{T^{\prime}}(a)=1$, $w_{T^{\prime}}(b)=w_{T^{\prime}}(c)=2$ and $w_{T^{\prime}}(d)=3$. We employ the notation $w_{T}$ for the weight function instead of $\varphi_{T}$ to indicate the common situation of a tree with positive real weights.
Step: $n=m=0$
This step is trivial since we only have pairs between leaves, like $(a,a^{\prime})$, which have trivial subtrees and thus $d_{E}(sub_{T}(a),sub_{T^{\prime}}(a^{\prime}))=0$.
Step: $n=m=1$
The points $x\in V_{T}$ with $\operatorname*{lvl}_{T}(x)\leq 1$ are $\{a,b,c,d\}$ and the points $y\in V_{T^{\prime}}$ with $\operatorname*{lvl}_{T^{\prime}}(y)\leq 1$ are $\{a^{\prime},b^{\prime},c^{\prime},d^{\prime}\}$. Thus the pairs $(x,y)$ which are considered are: $(d,d^{\prime})$, $(d,a^{\prime})$, $(d,b^{\prime})$, $(d,c^{\prime})$ and $(a,d^{\prime})$, $(b,d^{\prime})$, $(c,d^{\prime})$. The pairs between leaves, like $(a,a^{\prime})$ have already been considered.
Pair: $(d,d^{\prime})$
Let $T_{d}=sub_{T}(d)$ and $T_{d^{\prime}}=sub_{T^{\prime}}(d^{\prime})$. The set of internal vertices are respectively $E_{T_{d}}=\{a,b\}$ and $E_{T_{d^{\prime}}}=\{a^{\prime},b^{\prime}\}$. For each vertex $v<\text{root}$ in each subtree, where “root” stands for $d$ or $d^{\prime}$, roots of $T_{d}$ and $T_{d^{\prime}}$ respectively, we have $\zeta_{v}=\{v_{0}=v,v_{1}=\text{root}\}$. Thus, the binary variables we need to consider, are the following: $\delta_{0,0}^{a,a^{\prime}}$, $\delta_{0,0}^{a,b^{\prime}}$, $\delta_{0,0}^{b,a^{\prime}}$ and $\delta_{0,0}^{b,b^{\prime}}$. The quantities $\Delta_{i,j}^{v,w}$ are given by: $\Delta_{0,0}^{a,a^{\prime}}=0$, $\Delta_{0,0}^{a,b^{\prime}}=1$, $\Delta_{0,0}^{b,a^{\prime}}=0$ and $\Delta_{0,0}^{b,b^{\prime}}=1$. Thus:
$F^{C}(\delta)=0\cdot\delta_{0,0}^{a,a^{\prime}}+\delta_{0,0}^{a,b^{\prime}}+0\cdot\delta_{0,0}^{b,a^{\prime}}+\delta_{0,0}^{b,b^{\prime}}$
While:
$F^{D}(\delta)=(1-\delta_{0,0}^{a,a^{\prime}}-\delta_{0,0}^{a,b^{\prime}})\cdot 1+(1-\delta_{0,0}^{b,a^{\prime}}-\delta_{0,0}^{b,b^{\prime}})\cdot 1+(1-\delta_{0,0}^{a,a^{\prime}}-\delta_{0,0}^{b,a^{\prime}})\cdot 1+(1-\delta_{0,0}^{a,b^{\prime}}-\delta_{0,0}^{b,b^{\prime}})\cdot 2$
and:
$F^{-}(\delta)=(\delta_{0,0}^{a,a^{\prime}}+\delta_{0,0}^{a,b^{\prime}})\cdot 0+(\delta_{0,0}^{b,a^{\prime}}+\delta_{0,0}^{b,b^{\prime}})\cdot 0+(\delta_{0,0}^{a,a^{\prime}}+\delta_{0,0}^{b,a^{\prime}})\cdot 0+(\delta_{0,0}^{a,b^{\prime}}+\delta_{0,0}^{b,b^{\prime}})\cdot 0$
and:
$F^{S}(\delta)=\delta_{0,0}^{a,a^{\prime}}\cdot 0+\delta_{0,0}^{a,b^{\prime}}\cdot 0+\delta_{0,0}^{b,a^{\prime}}\cdot 0+\delta_{0,0}^{b,b^{\prime}}\cdot 0$
Lastly the constraints are:
$\delta_{0,0}^{a,a^{\prime}}+\delta_{0,0}^{a,b^{\prime}}\leq 1;\text{ }\delta_{0,0}^{b,a^{\prime}}+\delta_{0,0}^{b,b^{\prime}}\leq 1;\text{ }\delta_{0,0}^{a,a^{\prime}}+\delta_{0,0}^{b,a^{\prime}}\leq 1;\text{ }\delta_{0,0}^{a,b^{\prime}}+\delta_{0,0}^{b,b^{\prime}}\leq 1$
A solution is given by $\delta_{0,0}^{a,a^{\prime}}=\delta_{0,0}^{b,b^{\prime}}=1$ and $\delta_{0,0}^{a,b^{\prime}}=\delta_{0,0}^{b,a^{\prime}}=0$, which entails $F^{C}(\delta)=1$, $F^{D}(\delta)=0$, $F^{-}(\delta)=0$ and $F^{S}(\delta)=0$ and $d_{E}(T_{d},T_{d^{\prime}})=1$.
Pair: $(d,a^{\prime})$
Obviously: $d_{E}(sub_{T}(d),sub_{T^{\prime}}(a^{\prime}))=\mid\mid sub_{T}(d)\mid\mid$. All the pairs featuring a leaf and an internal vertex (that is, a vertex which is not a leaf), such as $(d,b^{\prime})$, $(a,d^{\prime})$ etc. behave similarly.
Step: $n=m=2$
The points $x\in V_{T}$ with $\operatorname*{lvl}_{T}(x)\leq 2$ are $\{a,b,c,d,r_{T}\}$ and the points $y\in V_{T^{\prime}}$ with $\operatorname*{lvl}_{T^{\prime}}(y)\leq 2$ are $\{a^{\prime},b^{\prime},c^{\prime},d^{\prime},r_{T^{\prime}}\}$. Thus the pairs $(x,y)$ which are considered are $(d,r_{T^{\prime}})$, $(r_{T},d^{\prime})$, $(r_{T},r_{T^{\prime}})$ and then the trivial ones: $(r_{T},a^{\prime})$, $(r_{T},b^{\prime})$, $(r_{T},c^{\prime})$ and $(a,r_{T^{\prime}})$, $(b,r_{T^{\prime}})$, $(c,r_{T^{\prime}})$. Some pairs have already been considered and thus are not repeated.
Pair: $(d,r_{T^{\prime}})$
Let $T_{d}=sub_{T}(d)$ and $T^{\prime}=sub_{T^{\prime}}(r_{T^{\prime}})$. The set of internal vertices are respectively $E_{T_{d}}=\{a,b\}$ and $E_{T_{d^{\prime}}}=\{a^{\prime},b^{\prime},c^{\prime},d^{\prime}\}$. Thus, the binary variables we need to consider, are the following: $\delta_{0,0}^{a,a^{\prime}}$, $\delta_{0,1}^{a,a^{\prime}}$, $\delta_{0,0}^{a,b^{\prime}}$, $\delta_{0,1}^{a,b^{\prime}}$, $\delta_{0,0}^{a,c^{\prime}}$, $\delta_{0,0}^{a,d^{\prime}}$, $\delta_{0,0}^{b,a^{\prime}}$, $\delta_{0,1}^{b,a^{\prime}}$, $\delta_{0,0}^{b,b^{\prime}}$, $\delta_{0,1}^{b,b^{\prime}}$ $\delta_{0,0}^{b,c^{\prime}}$, and $\delta_{0,0}^{b,d^{\prime}}$.
The quantities $\Delta_{i,j}^{v,w}$ are given by: $\Delta_{0,0}^{a,a^{\prime}}=0$, $\Delta_{0,1}^{a,a^{\prime}}=3$, $\Delta_{0,0}^{a,b^{\prime}}=1$, $\Delta_{0,1}^{a,b^{\prime}}=4$, $\Delta_{0,0}^{a,c^{\prime}}=1$, $\Delta_{0,0}^{a,d^{\prime}}=2$, $\Delta_{0,0}^{b,a^{\prime}}=0$, $\Delta_{0,1}^{b,a^{\prime}}=3$, $\Delta_{0,0}^{b,b^{\prime}}=1$, $\Delta_{0,1}^{b,b^{\prime}}=4$, $\Delta_{0,0}^{b,c^{\prime}}=1$ and $\Delta_{0,0}^{b,d^{\prime}}=2$. The function $F^{C}(\delta)$ is easily obtained by summing over $\delta_{i,j}^{v,w}\cdot\Delta_{i,j}^{v,w}$.
While:
$F^{D}(\delta)=(1-\delta_{0,0}^{a,a^{\prime}}-\delta_{0,1}^{a,a^{\prime}}-\delta_{0,0}^{a,b^{\prime}}-\delta_{0,1}^{a,b^{\prime}}-\delta_{0,0}^{a,c^{\prime}}-\delta_{0,0}^{a,d^{\prime}})\cdot 1+\ldots+(1-\delta_{0,0}^{a,d^{\prime}}-\delta_{0,0}^{b,d^{\prime}})\cdot 3$
and:
$F^{-}(\delta)=(\delta_{0,0}^{a,a^{\prime}}+\delta_{0,1}^{a,a^{\prime}}+\delta_{0,0}^{a,b^{\prime}}+\delta_{0,1}^{a,b^{\prime}}+\delta_{0,0}^{a,c^{\prime}}+\delta_{0,0}^{a,d^{\prime}})\cdot 0+\ldots+(\delta_{0,0}^{a,d^{\prime}}+\delta_{0,0}^{b,d^{\prime}})\cdot 3$
and:
$F^{S}(\delta)=(\delta_{0,0}^{a,a^{\prime}}+\delta_{0,1}^{a,a^{\prime}})\cdot 0+(\delta_{0,0}^{a,b^{\prime}}+\delta_{0,1}^{a,b^{\prime}})\cdot 0+\ldots+\delta_{0,0}^{a,d^{\prime}}\cdot 3+\delta_{0,0}^{b,d^{\prime}}\cdot 3$
Lastly the constraints are:
$\delta_{0,0}^{a,a^{\prime}}+\delta_{0,1}^{a,a^{\prime}}+\delta_{0,0}^{a,b^{\prime}}+\delta_{0,1}^{a,b^{\prime}}+\delta_{0,0}^{a,c^{\prime}}+\delta_{0,0}^{a,d^{\prime}}\leq 1$
$\delta_{0,0}^{b,a^{\prime}}+\delta_{0,1}^{b,a^{\prime}}+\delta_{0,0}^{b,b^{\prime}}+\delta_{0,1}^{b,b^{\prime}}+\delta_{0,0}^{b,c^{\prime}}+\delta_{0,0}^{b,d^{\prime}}\leq 1$
$\delta_{0,0}^{a,a^{\prime}}+\delta_{0,1}^{a,a^{\prime}}+\delta_{0,0}^{b,a^{\prime}}+\delta_{0,1}^{b,a^{\prime}}+\delta_{0,0}^{a,d^{\prime}}+\delta_{0,0}^{b,d^{\prime}}\leq 1$
$\delta_{0,0}^{a,b^{\prime}}+\delta_{0,1}^{a,b^{\prime}}+\delta_{0,0}^{b,b^{\prime}}+\delta_{0,1}^{b,b^{\prime}}+\delta_{0,0}^{a,d^{\prime}}+\delta_{0,0}^{b,d^{\prime}}\leq 1$
$\delta_{0,0}^{a,c^{\prime}}+\delta_{0,0}^{b,c^{\prime}}\leq 1$
In this case there are many minimizing solutions. One is given by: $\delta_{0,1}^{a,a^{\prime}}=\delta_{0,0}^{b,c^{\prime}}=1$ and all other variables equal to 00. This value of $\delta$ is feasible since the variables $\delta_{0,1}^{a,a^{\prime}}$ and $\delta_{0,0}^{b,c^{\prime}}$ never appear in the same constraint. This value of $\delta$ entails $F^{C}(\delta)=3+1$, $F^{D}(\delta)=2$, $F^{-}(\delta)=0$ and $F^{S}(\delta)=0$, and thus $d_{E}(T_{d},T^{\prime})=6$.
Another solution can be obtained with: $\delta_{0,0}^{a,d^{\prime}}=\delta_{0,0}^{b,c^{\prime}}=1$ and all other variables equal to 00. Also this value of $\delta$ is feasible since the variables $\delta_{0,0}^{a,d^{\prime}}$ and $\delta_{0,0}^{b,c^{\prime}}$ never appear in the same constraint. This value of $\delta$ entails $F^{C}(\delta)=2+1$, $F^{D}(\delta)=w_{T^{\prime}}(a^{\prime})+w_{T^{\prime}}(b^{\prime})=1+2$, $F^{-}(\delta)=\mid\mid sub_{T^{\prime}}(d^{\prime})\mid\mid=3$ and $F^{S}(\delta)=d_{E}(sub_{T}(a),sub_{T^{\prime}}(d^{\prime}))=\mid\mid sub_{T^{\prime}}(d^{\prime})\mid\mid=3$, and thus $d_{E}(T_{d},T^{\prime})=3+3-3+3=6$.
Pair: $(r_{T},d^{\prime})$
This and the other pairs are left to the reader.
B Proofs
Proof of 2.
We indicate with $\text{Lip}(\mathbb{R}^{2}_{\Delta})$ the set of functions $k:\mathbb{R}^{2}_{\Delta}\rightarrow\mathbb{R}$ which are Lipschitz on $\mathbb{R}^{2}_{\Delta}$ endued with the $1$-norm and such that $k(\Delta)=0$.
By Bubenik and Elchesen (2022) we have:
$W_{1}([D],[D^{\prime}])=W_{1}(\sum_{x\in[D]}\{x\},\sum_{y\in[D^{\prime}]}\{y\})=\sup_{k\in\text{Lip}(\mathbb{R}^{2}_{\Delta})}(\sum_{x\in[D]}k(x)-\sum_{y\in[D^{\prime}]}k(y)).$
At this point $(P3)$ and $(P4)$ follow easily by, respectively:
  • replacing $[D^{\prime}]$ with the empty diagram and noting:
    $\sup_{k\in\text{Lip}(\mathbb{R}^{2}_{\Delta})}(\sum_{x\in[D]}k(x))=\sum_{x\in[D]}\min_{y\in\Delta}\parallel x-y\parallel_{1};$
  • observing that when we have $W_{1}([D+D^{\prime\prime}],[D^{\prime}+D^{\prime\prime}])$ the contributions by the elements in $[D^{\prime\prime}]$ when computing:
    $\sum_{x\in[D+D^{\prime\prime}]}k(x)-\sum_{y\in[D^{\prime}+D^{\prime\prime}]}k(y)$
    appear in both sums, erasing themselves out.
$\blacksquare$
Proof of 5.
For the construction of the merge tree associated to a persistent set see Pegoraro (2024b).
Given a persistent set $S$ as in the hypotheses of the statement. To simplify the notation we make a slight abuse of notation and call $S$ also the restriction of $S$ to its critical set $C$.
Let $D_{S}$ be its display poset and $(T_{S},f_{S}:V\rightarrow\mathbb{R})$ with $T_{S}=(V,E)$ its merge tree. By construction $V-\{r_{T_{S}}\}$ with $r_{T_{S}}$ being the root of the merge tree and $f_{S}(r_{T_{S}})=+\infty$ can be sent into $D_{S}$: every vertex $v$ in $V-\{r_{T_{S}}\}$ is identified by a critical value $t_{i}=f_{S}(v)$ and an element $v\in S(t_{i})$. Thus we have $V-\{r_{T_{S}}\}\hookrightarrow D_{S}$. Moreover, this embedding is an embedding of posets if we consider $T_{S}$ with the relation generated by $child<parent$.
Consider now $(t_{j},p)\in D_{S}-V-\{r_{T_{S}}\}$: this happens if there are $t_{i}<t_{j}$ and $t_{k}>t_{j}$ such that for every $\varepsilon_{i}>0$ such that $t_{j}-\varepsilon_{i}>t_{i}$ and for every $\varepsilon_{k}>0$ such that $t_{j}+\varepsilon_{k}<t_{k}$, we have #$\#S(t_{j}\leq t_{j}+\varepsilon_{k})^{-1}(S(t_{j}\leq t_{j}+\varepsilon_{k})(p))=1$ and #$\#S(t_{j}-\varepsilon_{i}\leq t_{j})^{-1}(p)=1$. Moreover we can choose $t_{j}$ and $t_{k}$ so that:
  • $\max\{(t,s)\in D_{S}\mid(t,s)<(t_{j},p)\}=(t_{i},S(t_{i}\leq t_{j})^{-1}(p))$;
  • $\min\{(t,s)\in D_{S}\mid(t,s)>(t_{j},p)\}=(t_{k},S(t_{i}\leq t_{k})(p))$
This implies that for any edge $e$ connecting $(t,s)\rightarrow(t_{j},p)$ in the DAG representing $D_{S}$, we have that if $(t,s)\neq(t_{i},S(t_{i}\leq t_{j})^{-1}(p))$, then we also have an edge $(t,s)\rightarrow(t_{i},S(t_{i}\leq t_{j})^{-1}(p))$. But this implies that we can remove $e$ from the DAG without changing its transitive closure. That is, $e$ does not appear in $\mathcal{G}(D_{S})$. The same for any edge $(t_{j},p)\rightarrow(t,s)$ with $(t,s)\neq(t_{k},S(t_{i}\leq t_{k})(p))$. Thus $(t_{j},p)$ induces a degree $2$ vertex in $\mathcal{G}(D_{S})$.
To conclude the proof we just need to prove that $\mathcal{G}(D_{S})$ is a tree. In fact, if this is the case, then we have a poset map $V-\{r_{T}\}\hookrightarrow V_{\mathcal{G}(D_{S})}$ so that all the vertices not in the image of this map are of degree $2$. I.e. the map is an isomorphism of posets onto the image, which is the poset obtained by removing all degree $2$ vertices from $\mathcal{G}(D_{S})$.
We know that $\mathcal{G}(D_{S})$ is a DAG which has, by construction one maximal element. The only thing we need to prove is that for every $x=(t_{i},p)\in V_{\mathcal{G}(D_{S})}$ we only have one directed edge of the form $e=(x,y)$. But this is easily seen: $y=S(t_{i}\leq t_{i+1})(p)$. By functoriality, for every other $z=(t,s)\geq z$, we also have $z\geq y$. Thus we are done.
$\blacksquare$
Proof of 1.
Any degree $2$ vertex which is not ghosted is paired with another degree $2$ vertex. Ghosting both of them does not increase the cost of the mapping.
$\blacksquare$
Proof of 2.
To lighten the notation we use the following symbols:
  • the edit induced by $(v,\mathfrak{D})$ is called $v_{d}$ and $v^{-1}_{d}$ stands for $(\mathfrak{D},v)$.
  • the edit induced by $(v,\mathfrak{G})$ is called $v_{g}$ and $v^{-1}_{g}$ stands for $(\mathfrak{G},v)$.
  • the edit induced by $(v,v^{\prime})$ is called $v_{\varphi,\varphi^{\prime}}$ with $\varphi$ being the original weight function, and $\varphi^{\prime}$ the weight function after the shrinking.
We know that the set of finite edit paths between two dendrograms is nonempty.
Suppose that $\gamma$ is a finite edit path. This means that $\gamma$ is the composition of a finite set of edits. We indicate such ordered composition with $\gamma=\prod_{i=0}^{N}e_{i}$ with $e_{i}$ edit operation. We would like to change the order of the edit operations without raising the cost and changing the endpoints of the edit path. This is not always possible. However we can work it around in the useful cases using properties (P1)-(P4). In particular, we would like to know when we can commute a generic edit $e_{i}$ in the following situations:
  • $v_{d}\circ e_{i}$ and $e_{i}\circ v^{-1}_{d}$
  • $v_{g}\circ e_{i}$ and $e_{i}\circ v^{-1}_{g}$.
Moreover we want to reduce the edit path to max one edit for any vertex of $T$ and $T^{\prime}$.
We divide the upcoming part of the proof in subsections, each devoted to different combinations of edits.
Inverse Operations
We point out the following fact, which will be used often in the proof. Suppose we have a pair of edits $e_{i}\circ e_{j}$ that can be replaced with the pair $e^{\prime}_{i}\circ e^{\prime}_{j}$. Now consider $e^{-1}_{j}\circ e^{-1}_{i}$:
$e^{-1}_{j}\circ e^{-1}_{i}=(e_{i}\circ e_{j})^{-1}=(e^{\prime}_{i}\circ e^{\prime}_{j})^{-1}=e^{\prime-1}_{j}\circ e^{\prime-1}_{i}.$ (9)
$v_{d}$ and $v_{d}^{-1}$
When we delete or insert one vertex, we are modifying the tree structure at the level of its parent and its children. Therefore, we are only taking into account operations involving the parent of the considered vertex, on the vertex himself or on the children of the deleted/inserted vertex.
  • $v_{d}\circ v^{\prime}_{g}$: if $v$ is a child of $v^{\prime}$, can be safely replaced with $v_{d}\circ v^{\prime}_{d}$. Instead of ghosting the parent and then deleting the whole edge, we can delete both edges one by one; conserving the length of the path (P3). If, instead, $v$ is parent of $v^{\prime}$, we can safely commute the operations.
  • $v_{d}\circ v_{g}^{-1}$: the two edits can be replaced with $v^{\prime}_{\varphi,\varphi^{\prime}}$ with $v^{\prime}$ parent of $v$ (after the insertion) and $\varphi^{\prime}$ properly defined not to raise the cost of the path. In fact we are inserting $v$ on an edge and then deleting it. This can obviously be achieved by shrinking the original edge (without changing the path length - (P4)).
  • $v_{d}\circ v^{\prime-1}_{g}$: if $v^{\prime}$ is parent of $v$ after the splitting, we can again replace the two edits with shrinking: instead of inserting a point in an edge, and deleting then the edge below, we can directly shrink the original edge (P4). If, instead, $v^{\prime}$ is inserted below $v$, we have the same situation, but seen from the point of view of the vertex which is the child of $v^{\prime}$ after the splitting.
  • $v_{d}\circ v_{\varphi,\varphi^{\prime}}$: the two edits can be replaced by $v_{d}$ potentially diminishing the length of the path, but surely not raising it (P1).
  • $v^{\prime}_{g}\circ v_{d}^{-1}$: if $v^{\prime}$ is the parent of $v$, this edit can be replaced with just $v^{\prime}_{\varphi,\varphi^{\prime}}$ with appropriate weights: we are inserting an edge under a vertex which (in this case) becomes of degree two and is ghosted. We can directly modify the edge without changing the length of the path (P4). If, instead, $v^{\prime}$ becomes a child of $v$ after the insertion, we can simply shrink $v^{\prime}$ to obtain the same result without raising the cost (P4).
  • $v^{\prime-1}_{g}\circ v_{d}^{-1}$: if $v^{\prime}$ is on the edge inserted with $v_{d}^{-1}$, we cannot commute these two edits, but we can replace them with two insertions: instead of inserting an edge and then splitting it, we can directly insert two smaller edges; without changing the cost of the path (P3). In all other situations, we can commute the two operations.
  • $v_{\varphi,\varphi^{\prime}}\circ v_{d}^{-1}$: the two edits can be replaced with an insertion directly with weight $\varphi^{\prime}$, possibly shortening the path (P1).
  • $v^{\prime-1}_{d}\circ v_{d}$: suppose that the parent of $v^{\prime}$, after its insertion, is the parent of $v$ before we apply both edits. If the children of $v^{\prime}$ are different from the children of $v$, this operations do not commute. If the children are the same, they can be changed with a shrinking of $v$, reducing the length of the path by at most $cost(v^{\prime-1}_{d})+cost(v_{d})$ (P1).
$v_{g}$ and $v_{g}^{-1}$
Like in the previous case, we only take into account transformations concerning the parent and the child of the added/ghosted degree two vertex.
  • $v_{g}\circ v^{\prime}_{g}$: always commute (P2).
  • $v_{g}\circ v^{\prime-1}_{g}$: always commute, provided that we define carefully the splitting $v^{\prime-1}_{g}$ (P2).
  • $v_{g}\circ v_{\varphi,\varphi^{\prime}}$: with such operations, we are shrinking a vertex before ghosting it. However, we can achieve the same result, without increasing the path length, by ghosting the vertex at first, and then shrinking its child (P1)-(P4).
  • $v^{\prime}_{\varphi,\varphi^{\prime}}\circ v_{g}^{-1}$: if either $v^{\prime}=v$, or $v$ is the parent of $v^{\prime}$, we can replace those edits with an appropriate shrinking of the vertex which is the child of $v$ after the splitting and an appropriate insertion of $v^{\prime}$, without changing the length of the path (P3)-(P4).
  • $v_{g}\circ v^{\prime}_{d}$: if $v$ is the parent of $v^{\prime}$, such edits do not commute and can’t be replaced by a similar operation which inverts ghosting and deletion.
$v_{w,w^{\prime}}$
  • $v_{\varphi^{\prime},\varphi^{\prime\prime}}\circ v_{\varphi,\varphi^{\prime}}$: those edit can be replaced by $v_{\varphi,\varphi^{\prime\prime}}$ which is either conserving or shortening the path (P1).
  • $v_{\varphi,\varphi^{\prime}}^{-1}$: we can proceed as in the previous point, since: $v_{\varphi,\varphi^{\prime}}^{-1}=v_{\varphi^{\prime},\varphi}$.
Thanks to these properties, we can take a given path $\gamma=\prod_{i=0,...,N}e_{i}$ and modify the edit operations in order to obtain the following situation:
  • the first operations are all in the form $v_{d}$; this can be achieved because $v_{d}\circ-$ can be always rearranged, potentially by changing the path (as shown before) and shortening it. Of course, there can be only one deletion for each vertex of $T$;
  • then we have all the edits in the form $v_{g}$; since $v_{g}\circ-$ is exchangeable any time but when we have $v_{g}\circ v^{\prime}_{d}$, this is not a problem. Observe that all degree two vertices which were not deleted can be ghosted (at most one time);
  • in the same way we can put last all the paths in the form $v_{d}^{-1}$ and before them $v_{g}^{-1}$. All the new vertices appearing with the insertion of edges and the splitting of edges with degree two vertices are all nodes which remain in $T^{\prime}$ and which are not further edited;
  • in the middle we are left with the shrinking paths. Since we can substitute $v_{\varphi,\varphi^{\prime}}\circ v_{\varphi^{\prime},\varphi^{\prime\prime}}$ with $v_{\varphi,\varphi^{\prime\prime}}$, we can obtain at most one transformation on each vertex.
Thus, we can find at least one edit path of the form:
$\overline{\gamma}=(\gamma_{d}^{T^{\prime}})^{-1}\circ(\gamma_{g}^{T^{\prime}})^{-1}\gamma_{s}^{T}\circ\gamma_{g}^{T}\circ\gamma_{d}^{T}$
with:
  • $\gamma_{d}^{T}=\prod v_{d}$;
  • $\gamma_{g}^{T}=\prod v_{g}$;
  • $\gamma_{s}^{T}=\prod v_{\varphi,\varphi^{\prime}}$;
  • $(\gamma_{g}^{T^{\prime}})^{-1}=\prod v_{g}^{-1}$;
  • $(\gamma_{d}^{T^{\prime}})^{-1}=\prod v_{d}^{-1}$.
As for the paths $\gamma_{M}$ described in Section 4.2, all the permutations of the edits inside each $\gamma_{d}^{T},\gamma_{g}^{T},\ldots,$ does not change its endpoints and its cost. Thus, $\overline{\gamma}$ is such that $\gamma(T)=\overline{\gamma}(T)=T^{\prime}$ (up to isomorphism) and $cost(\overline{\gamma})\leq cost(\gamma)$.
The key point is that $\overline{\gamma}$ can be easily realized as a mapping made by the following elements:
  • $(v,\mathfrak{D})$, $\forall v_{d}\in\gamma_{d}^{T}$;
  • $(v,\mathfrak{G})$, $\forall v_{g}\in\gamma^{T}_{g}$;
  • $(v,v^{\prime})$, $\forall v_{\varphi,\varphi^{\prime}}\in\gamma_{s}^{T}$, where $v^{\prime}$ is the edge in $T^{\prime}$ (more precisely, in $\gamma_{g}^{T^{\prime}}\circ\gamma_{d}^{T^{\prime}}(T^{\prime})$) associated to the shrinking $v_{\varphi,\varphi^{\prime}}$;
  • $(\mathfrak{G},v)$, $\forall v^{-1}_{g}\in(\gamma_{g}^{T^{\prime}})^{-1}$;
  • $(\mathfrak{D},v)$, $\forall v^{-1}_{d}\in(\gamma_{d}^{T^{\prime}})^{-1}$.
$\blacksquare$
Proof of 6.
Condition (M2) coincide with condition (A2). Condition (M3) is clearly satisfied because of the antichain condition (A1). Consider a vertex $v\in E_{T}$. The only case in which $v$ is not edited is when $v<x$ with $x\in v_{T}\cap\pi_{T}(M^{*})$. However, in this case, $v$ does not appear in $\widetilde{T}$, and thus (M1) is satisfied. Moreover, after the deletions, all degree $2$ vertices are ghosted, and (M4) follows.
$\blacksquare$
Proof of 3.
Let $M\in M_{2}(T,T^{\prime})$ such that $d_{E}(T,T^{\prime})=cost(M)$.
We note that $parent>child$ induces a partial order relation also on the pairs given by paired points in $M$: $(x,y)>(v,w)$ if $x>v$ and $y>w$. In fact, by property (M3), $x>v$ if and only if $y>w$. So we can select $(x_{i},y_{i})$, the maxima with respect to this partial order relation . Thus, we obtain $(x_{0},y_{0})$,…,$(x_{n},y_{n})$ which form an antichain (both in $V_{T}$ and $V_{T^{\prime}}$).
Clearly $M^{*}=\{(x_{0},y_{0})$,…,$(x_{n},y_{n})\}\in\mathcal{C}^{*}(T,T^{\prime})$. Now we build $\alpha(M^{*})$ and compare the cost of its edits with the ones in $M$. Let $\bar{x}=LCA(x_{i},x_{j})$. Since $\bar{x}>x_{i},x_{j}$, it is not paired in $M$. Since $x_{i}$ and $x_{j}$ are paired, $\bar{x}$ cannot be ghosted, so it is deleted in $M$. Any point $x$ above $\bar{x}$ is deleted for the same reasons. So the edits above $\bar{x}$ are shared between $\alpha(M^{*})$ and $M$.
In $\alpha(M^{*})$ we ghost any point between $\bar{x}$ and $x_{i}$ (and the same for $x_{j}$) and this is not certain to happen in $M$ (some points could be deleted). Nevertheless, even in the worst case, these ghostings are guaranteed not to increase the distance. For instance, suppose $x_{i}<x<\bar{x}$ is deleted in $M$ and ghosted by $\alpha(M^{*})$, then:
$d(x_{i}\odot x,y_{i})\leq d(x_{i}\odot x,y_{i}\odot x)+d(y_{i}\odot x,y_{i})=d(x_{i},y_{i})+d(x,0)$
by properties (P1)-(P4). Since $\alpha(M^{*})\in M_{2}(\widetilde{T},\widetilde{T}^{\prime})$ by 6, we have:
$\sum_{(x,y)\in M^{*}}d_{E}(sub_{T}(x),sub_{T^{\prime}}(y))+cost(\alpha(M^{*}))\leq cost(M)$
Now we prove the other inequality.
Consider $M^{*}$ which realizes the minimum of the right side of Equation (3), and $M_{i}$ which realizes $d_{E}(sub(x_{i}),sub(y_{i}))$ with $(x_{i},y_{i})\in M^{*}$. We build a mapping $M$ collecting edits in the following way: for every $x^{\prime}\in E_{T}$ if $x^{\prime}\in sub(x_{i})$, we take the edit associated to it from $M_{i}$, otherwise we know that it is edited by $\alpha(M^{*})$, and we take it from there; the set of these assignments gives $M\in M_{2}(T,T^{\prime})$ whose cost is exactly $\sum_{(x_{i},y_{i})\in M^{*}}cost(M_{i})+cost(\alpha(M^{*}))$. This gives the second inequality.
$\blacksquare$
Proof of 7.
Having fixed a leaf $l$, the constraint in Equation (5) allows for at most one path $\zeta_{v}^{v_{i+1}}\subset\zeta_{l}$ to be kept after the edits induced by all the variables equal to $1$. Moreover if $(v^{\prime\prime},i)\in\Phi(v)\cap\Phi(v^{\prime})$, then $v=v^{\prime\prime}_{i}=v^{\prime}$. Thus, variables are added at most one time in Equation 5 and Equation 6. Which means that for any $a\in V_{T_{x}}$, we are forcing that the vertex $a$ can be an internal vertex or lower extreme of at most one path $\zeta_{v}^{v_{i+1}}$ such that $\delta_{i,j}^{v,w}=1$. In other words if two “kept” paths $\zeta_{v}^{v_{i+1}}$ and $\zeta_{v^{\prime}}^{v^{\prime}_{i^{\prime}+1}}$ (i.e. with $\delta_{i,j}^{v,w}=\delta_{i^{\prime},j}^{v^{\prime},w}=1$) intersect each other, it means that they just share the upper extreme $v_{i+1}=v^{\prime}_{i^{\prime}+1}$. These facts together imply that (if the constraints are satisfied) the edits induced on $T_{x}$ by $\delta^{v,w}_{i,j}=1$ and $\delta^{v^{\prime},w^{\prime}}_{i^{\prime},j^{\prime}}=1$ are always compatible: if $v^{\prime\prime}\in V_{sub_{T}(v_{i})}$ then it is not touched by (the edits induced by) $\delta^{v^{\prime},w^{\prime}}_{i^{\prime},j^{\prime}}=1$ (and the same exchanging the role of $v^{\prime}$ and $v$), if $v^{\prime\prime}$ is equal or above $v_{i+1}$ and/or $v^{\prime}_{i^{\prime}+1}$, then it is deleted in any case. Lastly, by noticing that if $\delta^{v,w}_{i,j}\in\Phi(v^{\prime})$ then $\delta^{v,w^{\prime}}_{i,j^{\prime}}\in\Phi(v^{\prime})$ for all other possible $w^{\prime}$ and $j^{\prime}$, we see that every path $\zeta_{v}^{v_{i}}$ is paired with at most one path $\zeta_{w}^{w_{j}}$, and viceversa.
As a consequence, for any vertex $v^{\prime}$ in any of the tree structures, at most one point on the path $\zeta_{v^{\prime}}$ is paired in $M^{*}$, guaranteeing the antichain condition. Moreover, any point of $T_{x}$ which is in $\pi_{T_{x}}(M^{*})$ is assigned to one and only one point of $T^{\prime}_{y}$ and viceversa. The edits induced by $\delta=1$ clearly satisfy properties (M2)-(M4). Passing to $\widetilde{T}_{x}$ and $\widetilde{T}_{y}$, also (M1) is satisfied.
$\blacksquare$